### v1

# DivCon: Divide and Conquer for Progressive Text-to-Image Generation

Abstract
Diffusion-driven text-to-image (T2I) generation has achieved remarkable advancements. To further improve T2I models’ capability in numerical and spatial reasoning, the layout is employed as an intermedium to bridge large language models and layout-based diffusion models. However, these methods still struggle with generating images from textural prompts with multiple objects and complicated spatial relationships. To tackle this challenge, we introduce a divide-and-conquer approach which decouples the T2I generation task into simple subtasks. Our approach divides the layout prediction stage into numerical & spatial reasoning and bounding box prediction. Then, the layout-to-image generation stage is conducted in an iterative manner to reconstruct objects from easy ones to difficult ones. We conduct experiments on the HRS and NSR-1K benchmarks and our approach outperforms previous state-of-the-art models with notable margins. In addition, visual results demonstrate that our approach significantly improves the controllability and consistency in generating multiple objects from complex textural prompts.
Keywords:
Text-to-Image Generation Large Language Models Diffusion Models
Figure 1: Layouts and images generated by our DivCon. DivCon enhances the ability of text-to-image diffusion models to understand complex numerical and spatial relationships in the text.
1 Introduction
Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects [30, 7, 3, 21, 29, 31]. However, these methods still suffer from limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships [2, 7, 14, 8, 35]. Recently, combining Large Language Models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ ability in numerical and spatial reasoning [15, 26]. LLMs have strong reasoning and visual planning ability to predict objects’ layouts from the text prompt, which provides more informative conditions for diffusion models to produce higher-quality images.
Despite the significant advancement, existing methods still suffer from limited accuracy when the text prompt contains multiple objects and complicated spatial relationships [2, 26]. Intuitively, the difficulty of synthesizing objects with different characteristics varies a lot. Specifically, LLMs encounter great challenges when conducting simultaneous reasoning and visual planning from complicated text prompts (e.g. "five mice are scurrying around a wooden table, each nibbling on one of the five ripe apples that are laid out on top."). Besides, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and details. However, previous methods generate all objects simultaneously without taking their differences into consideration. As a result, objects with higher difficulties cannot be well generated in the images (e.g. toaster and pizza, Fig. 2).
To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation phases. Our approach divides complicated tasks into multiple subtasks according to their difficulties and then address these subtasks in an iterative manner. During layout prediction, the number of objects and their spatial relationships are first reasoned from the text prompt. Then, visual planning is performed to predict object bounding boxes. During layout-to-image generation, the diffusion model is first executed to process the layout. Then, the consistency between the synthetic objects and the text prompt is calculated. Next, low-fidelity objects’ layouts are highlighted and fed to the diffusion model to encourage the network to focus on objects with higher difficulties. Extensive experiments have demonstrated that DivCon has significantly improved the quality of generated and achieves state-of-the-art performance on the HRS[2] and NSR-1K[15] benchmark datasets.
Our main contributions can be summarized as follows:
  • We propose DivCon, a divide-and-conquer approach for layout-based text-to-image generation by dividing complicated tasks into simple subtasks.
  • To obtain accurate layouts from the text prompt, we divide the layout prediction stage into numerical & spatial reasoning and bounding box planning.
  • To generate high-fidelity images from the layout, we divide layout-to-image generation into two iterations to synthesize objects with different levels of difficulty.
  • Experimental results show that our approach achieves significant performance gains, surpassing current state-of-the-art models.
Figure 2: A comparison of generation difficulty between "toaster" and "pizza". Current methods tend to generate pizza but fail to generate enough toasters or completely miss the category.
2 Related work
In this section, we review three main categories of text-to-image generation methods: text-to-image generation (T2I) models, layout-conditioned T2I models, and text-to-layout-to-image generation models (Fig. 3).
2.0.1 Text-to-Image (T2I) Generation Models
The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions [3, 17, 29, 21, 31, 38]. Milestone methods in this area are Generative Adversarial Networks (GANs) [21, 32], Variational Autoencoders (VAEs) [6, 12, 29, 38], and diffusion models [3, 19, 25, 28, 31]. GANs often struggle with training stability and mode collapse, whereas VAEs may sometimes produce overly smooth or blurry outputs due to their reconstruction loss. Recently, diffusion models are highly valued for their robustness and ability to generate detailed and realistic images. Particularly, Stable Diffusion[30], which utilizes open vocabulary encoder[27] for encoding text to embeddings as conditional inputs, has demonstrated a remarkable capability for open-world image synthesis, pushing the boundaries of generative models. Despite promising results, a text prompt is usually difficult to provide fine-grained control of the generated images.
Figure 3: Three main categories of text-to-image generation methods.
2.0.2 Layout-conditioned T2I Models
To improve the controlability of T2I models, prompts like layouts as employed as additional conditions for T2I models. Existing layout-conditioned generation models can be categorized into three branches. One branch of methods trains external networks on layout-images pairs [22, 39, 25, 1, 37]. For instance, GLIGEN incorporates a trainable gated self-attention layer that transforms bounding box information into conditional input for Stable Diffusion. Another branch of methods, including attention-refocusing [26], layout-predictor [34], direct-diffusion[24], layout-guidance[9] and BoxDiff [36], directly manipulate attention maps to optimize the sampling process. These methods generate images with multiple objects through a single forward pass of the diffusion model, resulting in limited fidelity for objects with complicated appearances or spatial relationships. To remedy this, the last branch of methods including Mixture-of-Diffusion[5], MultiDiffusion [4], and LLM-grounded Diffusion [23] perform denoising on each object (i.e., bounding box) separately before fusing the results. However, these methods introduce much higher computational costs. Moreover, the requirements of additional layout prompts hinder their practicality.
2.0.3 Text-to-Layout-to-Image Generation Models
Current LLMs have demonstrated strong numerical and spatial reasoning abilities in visual planning and layout generation. Inspired by this, several efforts are made to integrate LLMs to generate layouts from texts, and then utilize layout-conditioned diffusion models for image generation. Attention-refocusing [26] and LLM-grounded Diffusion [23] explore in-context learning of GPT-4 for layout prediction. LayoutGPT [15] searches similar samples from their database as input examples for the LLMs. LLM-layout-generator [10] leverages Chain-of-Thought (CoT) prompting [33] to improve the layout prediction accuracy. However, LLMs are expected to simultaneously perform numerical & spatial reasoning and visual planning of bounding boxes in these methods. Contrary to these methods, we propose a divide-and-conquer approach to conduct reasoning and visual planning separately, which improves the accuracy of layout predictions and substantially advances the prompt fidelity of generated images. In addition, during layout-to-image generation, the objects are synthesized in an iterative manner according to their difficulties, which achieves not only superior fidelity but also significant computational cost savings.
Figure 4: The proposed DivCon framework. The layout prediction stage is divided into numerical & spatial reasoning and bounding box planning. The layout-to-image generation is divided into two iterations to generate objects with different difficulty levels separately.
3 Method
3.1 Overview
As illustrated in Fig. 4, our framework comprises two stages. Given a text prompt, LLMs first predict layouts from it. Then, the resultant layout is employed as an additional condition to achieve text-to-image generation. More specifically, each stage is further divided into two steps to handle simple subtasks.
3.2 Text-to-Layout Prediction
Intuitively, to predict the layout from a text prompt, the number & spatial relations of the objects are first required to be reasoned, and then bounding boxes are generated to describe the coordinates & sizes of the objects. However, previous methods commonly conduct these two steps simultaneously without considering their inherent correlation, thereby hindering them from understanding complicated text prompts well. To remedy this, we introduce the idea of divide-and-conquer to improve the accuracy of predicted layouts. Specifically, leveraging the state-of-the-art large language model GPT-4 [], numerical and spatial reasoning is first performed, and then bounding boxes are predicted.
3.2.1 Numerical and Spatial Reasoning
Given a text prompt $C$ describing an image, an LLM is adopted to predict a set of strings $O=\{s_{j}|j=1,2,...,n\}$. If the prompt primarily focused on counting relationships among objects, each string $s_{j}$ includes both the name and quantity information of each object category $j$. If the prompt mainly describes spatial relationships, each string $s_{j}$ will contain the object name and its spatial position in the image. This method allows LLMs to conclude complex numerical and spatial relationships.
3.2.2 Bounding Box Prediction
After numerical and spatial reasoning, both the meta prompt $C$ and the prediction result $O$ are fed to the LLM again to predict a list of name-box pairs $P=\{p_{i}|i=1,2,...,n\}$. Each pair $p_{i}=\{o_{i}:(l_{i},t_{i},r_{i},b_{i})|i=1,2,...,n\}$ consists of the object name $o_{i}$ and its bounding box’s coordinates $(l_{i},t_{i},r_{i},b_{i})$. The in-context exemplars include a process of correcting the features of generated bounding boxes to ensure the quality of the output layouts.
3.3 Layout-to-Image Generation
After layout prediction, the resultant layout is employed as an additional condition to generate the image. As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy to hard in an iterative manner, as shown in Fig. 4. First, the layout is fed to the diffusion model to synthesize an image. Then, the consistency between the synthetic objects and the text prompt is calculated. Afterward, objects with low fidelity are highlighted and passed to the diffusion model for a second time, encouraging the diffusion model to focus on these difficult objects to achieve higher quality.
3.3.1 First-Round Generation
We employ a layout-to-image diffusion model with sampling optimization [26] as the generative model to generate an image conditioned on the text prompt and the predicted layout prompt. During the denoising process, random noise $z_{T}$ is fed to the diffusion model to gradually denoise it to an image $z_{0}$ after $T$ iterations.
3.3.2 Consistency Evaluation
With the generated image $x$, each object is cropped using the bounding boxes in the layout, which are denoted as $x_{i}\in\mathbb{R}^{4\times h\times w}$, where $h\times w$ represents the height and width of the bounding box. Then, the CLIP similarity $S_{x_{i},o_{i}}\in(-1,1)$ is employed to evaluate the consistency between the cropped patch and the object name $o_{i}$ in the text prompt. The higher the similarity is, the higher the fidelity of the corresponding object is. Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are "good enough”. In contrast, other objects with low similarities are hard samples that are more difficult to reconstruct (e.g. "sheep" and "horse" in Fig. 5). Next, a binary mask $M$ is calculated with bounding boxes of easy objects being set to 1 and the rest regions being set to 0.
Figure 5: Comparison of easy and hard samples in a generated image.
3.3.3 Second-Round Refinement
After consistency evaluation, different objects are of various levels of fidelity. Our objective is to preserve the “good enough” objects while reproducing the remaining ones to ease the difficulty of diffusion model. To this end, we retrieve the latent representation during the denoising process of the first-round generation. Specifically, $z_{0.7T}$ and $z_{0.3T}$ are first retrieved. Then, the binary mask $M$ is adopted to crop high-fidelity regions in $z_{0.7T}$ and $z_{0.3T}$. Next, the starting noise $z^{\prime}_{T}$ for the second-round diffusion is obtained as:
$z^{\prime}_{T}=z_{0.7T}\circ M+\mathcal{N}(\mu,\sigma^{2})\circ(1-M)$ (1)
where $\circ$ is element-wise multiplication and $\mathcal{N}(\mu,\sigma^{2})$ is random noise.
The second-round denoising process consists of three stages. First, the input noise $z^{\prime}_{T}$ is denoised for $0.3T$ iterations to obtain $z^{\prime}_{0.3T}$. Second, the high-fidelity regions in $z_{0.7T}$ are cropped and used to replace the corresponding regions in $z^{\prime}_{0.3T}$ and fed to the diffusion model for anthoer $0.4T$ denoising iterations. Note that, the text prompt and the full layout prompt are employed as conditions in these two stages. Moreover, the regions marked by $M$ is frozen to preserve these high-fidelity regions. During optimization, Cross-Attention Refocusing (CAR) and Self-Attention Refocusing (SAR) losses [26] are employed to update only regions of low fidelity, encouraging the model to focus on these hadr samples. Finally, all regions are optimized for the last $0.3T$ iterations to smoothly blend the high-fidelity regions with other regions.
4 Experiments
In this section, we provide an evaluation of our method and compare it with state-of-the-art T2I models. Ablation experiments are conducted to demonstrate the effectiveness of our divide-and-conquer approach in layout prediction and layout-to-image generation stages.
4.1 Experiment Setup
4.1.1 Benchmarks
To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS[2] and NSR-1K[15]. The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships. The NSR-1K dataset consists of 762 prompts for numerical relationships and 283 prompts for spatial relationships, all labeled with names, counts, bounding boxes, and ground truth images. We also utilize NSR-1K for quality evaluation. Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations. This elevates HRS as a more sophisticated and challenging benchmark.
4.1.2 Evaluation Metrics
We evaluate the images produced by our methods in terms of both quality and fidelity.
(1) Quality To assess the quality of generated images, we use FID as the metric[18]. Lower FID represents higher similarity between generated images and ground truths images. We report FID in NSR-1K [15] as each prompt has a corresponding ground truth image for FID calculation.
(2) Fidelity To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated. Particularly, YOLOv8 [20] is employed to obtain the bounding boxes of the objects in the generated images. In counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy [13, 16]. In spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills[11] and LayoutGPT [15]. Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship. Otherwise, they are denoted as a top-bottom relationship.
4.1.3 Baselines
Stable Diffusion [30], Attend-and-Excite[7], and Attention-Refocusing [26] are employed for comparison. Stable Diffusion and Attend-and-Excite directly generate images from the text prompt. Meanwhile, Attention-Refocusing employs layout as an intermedium for image generation.
Table 1: Comparison of our method with baselines on HRS benchmark. Text in bold shows the best results of each metric.
Methods
Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc.
Spatial
Layout Image
Acc.
Acc.
Stable Diffusion 2.1[30]
- - - - 73.97 53.93 62.38 16.20
7.96
Attend-and-Excite[7]
- - - - 76.57 55.94 64.65 18.37
9.23
Attention-Refocusing[26]
88.62 91.91 90.44 77.05 77.43 61.78 68.72 25.60
89.57
48.29
DivCon (Ours)
94.17 89.53 91.80 80.73 79.13 63.53 70.48 29.62
91.87
53.87
Table 2: Comparison of our method with baselines on NSR-1K benchmark. Text in bold shows the best results of each metric.
Methods
Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc.
Spatial
Layout Image
Acc.
Acc.
Stable Diffusion 2.1[30]
- - - - 78.94 75.42 77.41 33.99
22.97
Attend-and-Excite[7]
- - - - 79.21 78.62 78.91 39.51
29.61
Attention-Refocusing[26]
98.51 99.11 98.81 94.31 84.61 85.64 85.12 56.10
95.05
69.28
DivCon (Ours)
98.47 98.35 98.41 94.39 85.31 87.54 86.41 55.80
96.11
72.95
Figure 6: Qualitative comparison of numerical reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing, and our DivCon.
Figure 7: Qualitative comparison of spatial reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing and our DivCon.
Table 3: FID results achieved by different methods in NSR-1K.
Methods
FID ($\downarrow$)
Stable Diffusion 2.1[30]
20.77
Attention-Refocusing[26]
21.05
DivCon (Ours)
20.89
4.2 Evaluation Results
4.2.1 Quantitative Results
We present the results on the HRS benchmark [2] in Tab. 1. As we can see, DivCon outperforms previous T2I models by over 10% in numerical accuracy and over 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance. In addition, DivCon surpasses Attention Refocusing [26] on most metrics. Specifically, our layout prediction improves numerical accuracy by 3% and spatial accuracy by 2%, respectively. For image generation, the F1 score and accuracy achieve 2% and 4% gains, while the spatial accuracy is boosted by up to 6%. By dividing the T2I task into simple subtasks, our DivCon can better understand the text prompt to achieve higher fidelity.
The results on the NSR-1K benchmark are presented in Tab. 2. It can be observed that DivCon outperforms other T2I models by over 15% in numerical accuracy and over 45% in spatial accuracy. As compared to Attention Refocusing [26], our DivCon produces competitive results in terms of layout accuracy. Meanwhile, DivCon achieves approximately 2% superiority in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy.
Quantitative results in terms of image quality are shown in Tab. 3. As we can see, our DivCon produces compatible FID scores as compared to stable diffusion and outperforms attention-refocusing with notable margins. This demonstrates that our approach is able to preserve image quality while enhancing prompt fidelity.
4.2.2 Qualitative Results
Fig. 6 and Fig. 7 illustrate the qualitative comparison of DivCon and baselines. In all cases, our DivCon can accurately predict the object quantities and the spatial relations from the text prompts.
(1) Large Quantity of Objects. Given a prompt containing more than one object category (Fig. 6), Stable Diffusion [30] and Attend-and-Excite [7] often generate an insufficient quantity of objects, while Attention-Refocusing [26] sometimes omits one category of objects or generates an excessive number of objects. For instance, the prompt in Fig. 6(b) contains five spoons and four oranges. However, Stable Diffusion and Attend-and-Excite cannot synthesize sufficient spoons or oranges. Meanwhile, Attention-Refocusing generates an excess number of objects, with spoons appearing incomplete due to missing parts. As compared to these methods, our DivCon achieves superior numerical accuracy.
(2) Complicated Spatial Relation. Given a prompt containing complex spatial relationships among more than two object categories (Fig. 7, previous T2I models struggle to generate correct spatial relations. Meanwhile, Attention-Refocusing tends to produce distortion or blending artifacts around the boundaries of objects. In Fig. 7 (b), Stable Diffusion and Attend-and-Excite miss certain object categories, while Attention-Refocusing merges the color of the banana with the person. In contrast, our DivCon inference accurate object categories and their spatial relationships, exhibiting superior performance in these challenging cases.
Figure 8: Comparison of layouts generated from Attention-Refocusing and our DivCon.
Figure 9: Ablation study of our image generation.
Figure 10: Comparison of transmit one or two latent representations.
4.3 Ablation Study
In this sub-section, we conduct ablation experiments to evaluate the effectiveness of our divide-and-conquer approach in both layout prediction and image generation. Results are presented in Tab. 4 and Attention-Refocusing [26] is used as the baseline.
4.3.1 Divide-and-Conquer in Layout Prediction
To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step. As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 4% in both numerical and spatial accuracy. We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig. 8. Compared to Attention-Refocusing, layouts generated by our DivCon not only exhibit higher reasoning accuracy (Fig. 8 (a), (b)) but also are more organized with less overlap (Fig. 8 (c), (d)).
4.3.2 Divide-and-Conquer in Image Generation
To study the effectiveness of our divide-and-conquer approach in the image generation stage, we develop a variant to synthesize the image in a single forward pass. It can be observed from Tab. 4 that the divide-and-conquer approach facilitates our method to notable performance gains in terms of all metrics. We further compare the results produced by our method using prompts containing multiple objects in Fig. 9. As we can see, the divide-and-conquer approach facilitates our DivCon to better reconstruct the objects, especially for small-sized ones, while avoiding generating objects outside the designated bounding boxes. For example, in Fig. 9(a), DivCon accurately reconstructs the small objects. In Fig. 9(d), DivCon successfully avoids generating the dog outside boxes.
Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy.
4.3.3 Fractiles in Image Generation
During image generation of our DivCon, two fractiles (i.e., $z_{0.3T}$ and $z_{0.7T}$) are used to preserve “good enough” results in the first-round generation. We conduct ablation experiments to study these two fractiles. The second fractile $z_{0.7T}$ marks the end of optimization and is non-adjustable. We observe that transmitting only $z_{0.7T}$ to the next phase results in instability in image generation instability, specifically manifesting as noticeable boundary demarcations of bounding boxes in the images (Fig. 10).
We further compare the performance achieved by our model with different values for the first fractile. Experiment results are presented in Tab. 5. It can be observed that the model suffers a notable performance drop if only the second fractile is employed. With an additional fractile, our DivCon yields a notable improvement. Additionally, the selection of the first fractile’s value exhibits a marginal impact on the efficacy of the model, with the 30% threshold demonstrating optimal performance.
Table 4: Ablation study of our method’s image generation performance on HRS benchmark. "Layout w/DivCon": Layout prediction from LLMs with divide-and-conquer prompting. "Image w/DivCon": Layout-to-Image generation with divide-and-conquer strategy.
Layout w/DivCon Image w/DivCon
Numerical
Precision
Recall
F1
Acc.
Spatial
Acc.
77.43
61.78
68.72
25.60
48.29
78.26
62.61
69.57
29.46
52.74
78.18
62.36
69.38
27.83
53.17
79.13
63.53
70.48
29.62
53.87
Table 5: Ablation study of two fractiles used to preserve results in the first-round generation.
First Fractile Second Fractile
Numerical
Precision
Recall
F1
Acc.
Spatial
Acc.
76.82
60.69
67.80
24.31
48.53
$0.1T$
78.44
61.61
69.01
29.12
52.30
$0.3T$
79.13 63.53 70.48 29.62
53.87
$0.5T$
78.13
63.25
69.90
28.71
51.46
5 Limitations and Discussions
As illustrated in Fig. 11, our method still struggles with generating objects from certain pattern layouts. When the text prompt describes objects overlapping in space, the diffusion model may fail to generate all objects accurately. This limitation stems from the inherent capacity of layout-conditioned image generation models to handle overlapping boxes. Despite these challenges, our approach still demonstrates a reduced frequency of overlapping in layout prediction compared to the base model.
Figure 11: Comparison of transmit one or two latent representations.
6 Conclusion
In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into simple subtasks to achieve higher image generation fidelity. Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, respectively. Then, layout-to-image generation is achieved via an iterative manner to synthesize objects in an easy-to-hard fashion. Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.

### v2

@@
 Abstract
-Diffusion-driven text-to-image (T2I) generation has achieved remarkable advancements.
-To further improve T2I models’ capability in numerical and spatial reasoning, the layout is employed as an intermedium to bridge large language models and layout-based diffusion models.
-However, these methods still struggle with generating images from textural prompts with multiple objects and complicated spatial relationships.
-To tackle this challenge, we introduce a divide-and-conquer approach which decouples the T2I generation task into simple subtasks.
-Our approach divides the layout prediction stage into numerical & spatial reasoning and bounding box prediction.
-Then, the layout-to-image generation stage is conducted in an iterative manner to reconstruct objects from easy ones to difficult ones.
-We conduct experiments on the HRS and NSR-1K benchmarks and our approach outperforms previous state-of-the-art models with notable margins.
-In addition, visual results demonstrate that our approach significantly improves the controllability and consistency in generating multiple objects from complex textural prompts.
-Keywords:
-Text-to-Image Generation Large Language Models Diffusion Models
-Figure 1: Layouts and images generated by our DivCon. DivCon enhances the ability of text-to-image diffusion models to understand complex numerical and spatial relationships in the text.
-1 Introduction
-Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects [30, 7, 3, 21, 29, 31].
-However, these methods still suffer from limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships [2, 7, 14, 8, 35].
-Recently, combining Large Language Models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ ability in numerical and spatial reasoning [15, 26].
-LLMs have strong reasoning and visual planning ability to predict objects’ layouts from the text prompt, which provides more informative conditions for diffusion models to produce higher-quality images.
-Despite the significant advancement, existing methods still suffer from limited accuracy when the text prompt contains multiple objects and complicated spatial relationships [2, 26].
-Intuitively, the difficulty of synthesizing objects with different characteristics varies a lot.
-Specifically, LLMs encounter great challenges when conducting simultaneous reasoning and visual planning from complicated text prompts (e.g.
-"five mice are scurrying around a wooden table, each nibbling on one of the five ripe apples that are laid out on top.").
-Besides, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and details.
-However, previous methods generate all objects simultaneously without taking their differences into consideration.
-As a result, objects with higher difficulties cannot be well generated in the images (e.g.
-toaster and pizza, Fig.
+Diffusion-driven text-to-image (T2I) generation has achieved remarkable advancements in recent years.
+To further improve T2I models’ capability in numerical and spatial reasoning, layout is employed as an intermedium to bridge large language models and layout-based diffusion models.
+However, these methods still struggle with generating images from prompts with multiple objects and complicated spatial relationships.
+To tackle this challenge, we introduce a divide-and-conquer approach which decouples the generation task into multiple subtasks.
+First, the layout prediction stage is divided into numerical & spatial reasoning and bounding box prediction to better parse the text prompt to obtain accurate composition information.
+Second, the layout-to-image generation stage is divided into two steps to synthesize objects from easy ones to difficult ones.
+Experiments are conducted on the HRS and NSR-1K benchmarks and our method outperforms previous approaches with notable margins.
+In addition, visual results and user study demonstrate that our approach significantly improves the perceptual quality, especially when generating multiple objects from complex textural prompts.
+Codehttps://github.com/DivCon-gen/DivCon
+Introduction
+Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects (Rombach et al.
+2022; Chefer et al.
+2023; Balaji et al.
+2022; Kang et al.
+2023; Ramesh et al.
+2021; Saharia et al.
+2022).
+However, these methods still suffer limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships (Bakr et al.
+2023; Chefer et al.
+2023; Feng et al.
+2023a; Chen, Laina, and Vedaldi 2023a; Xiao et al.
+2023).
+Recently, combining large language models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ capability in numerical and spatial reasoning (Feng et al.
+2023b; Phung, Ge, and Huang 2024).
+These methods leverage the strong reasoning and visual planning ability of LLMs to predict objects’ layouts from the text prompt, which provides additional informative conditions for diffusion models to produce higher-quality images.
+Despite recent significant advancements, existing methods still suffer limited accuracy when the text prompt contains multiple objects and complicated spatial relationships (Bakr et al.
+2023; Phung, Ge, and Huang 2024).
+On the one hand, LLMs encounter great challenges when conducting simultaneous reasoning and visual planning from complicated text prompts (e.g.
+“five mice are scurrying around a wooden table, each nibbling on one of the five ripe apples that are laid out on top”).
+On the other hand, since the difficulty of synthesizing objects with different characteristics varies a lot, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and details.
+However, previous methods commonly conduct reasoning and visual planning together while generating all objects simultaneously without taking their differences into consideration.
+As a result, objects with higher difficulties cannot be well synthesized in the images (e.g.
+toaster and pizza in Fig.
 2).
-To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation phases.
-Our approach divides complicated tasks into multiple subtasks according to their difficulties and then address these subtasks in an iterative manner.
+Figure 1: Layouts and images generated by our DivCon. DivCon enhances the capability of text-to-image diffusion models to understand complex numerical and spatial relationships in the text.
+To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation stages.
 During layout prediction, the number of objects and their spatial relationships are first reasoned from the text prompt.
-Then, visual planning is performed to predict object bounding boxes.
-During layout-to-image generation, the diffusion model is first executed to process the layout.
-Then, the consistency between the synthetic objects and the text prompt is calculated.
-Next, low-fidelity objects’ layouts are highlighted and fed to the diffusion model to encourage the network to focus on objects with higher difficulties.
-Extensive experiments have demonstrated that DivCon has significantly improved the quality of generated and achieves state-of-the-art performance on the HRS[2] and NSR-1K[15] benchmark datasets.
+Then, visual planning is conducted to predict the bounding boxes of the objects.
+By decoupling layout prediction into reasoning and visual planning, more accurate composition information can be obtained.
+During layout-to-image generation, the diffusion model is first executed to process the layout to generate all objects.
+Then, the consistency between the resultant objects and the text prompt is calculated.
+With objects of low consistency being highlighted, the diffusion model is executed for another time to encourage the network to focus more on these difficult objects.
+By diving layout-to-image generation into two steps, objects with different levels of difficulties can be synthesized progressively for higher quality.
+Extensive experiments have demonstrated that our DivCon has significantly improved the quality of generated images and achieves state-of-the-art performance on the HRS(Bakr et al.
+2023) and NSR-1K(Feng et al.
+2023b) benchmark datasets.
 Our main contributions can be summarized as follows:
-  • We propose DivCon, a divide-and-conquer approach for layout-based text-to-image generation by dividing complicated tasks into simple subtasks.
-  • To obtain accurate layouts from the text prompt, we divide the layout prediction stage into numerical & spatial reasoning and bounding box planning.
-  • To generate high-fidelity images from the layout, we divide layout-to-image generation into two iterations to synthesize objects with different levels of difficulty.
+  • We propose DivCon, a divide-and-conquer approach for layout-based text-to-image generation by dividing the complicated task into multiple subtasks.
+  • To obtain accurate composition information from the text prompt, we divide the layout prediction stage into numerical & spatial reasoning and bounding box planning.
+  • To generate high-fidelity images from the layout, we divide layout-to-image generation into two steps to synthesize objects with different levels of difficulty.
   • Experimental results show that our approach achieves significant performance gains, surpassing current state-of-the-art models.
-Figure 2: A comparison of generation difficulty between "toaster" and "pizza". Current methods tend to generate pizza but fail to generate enough toasters or completely miss the category.
-2 Related work
-In this section, we review three main categories of text-to-image generation methods: text-to-image generation (T2I) models, layout-conditioned T2I models, and text-to-layout-to-image generation models (Fig.
-3).
-2.0.1 Text-to-Image (T2I) Generation Models
-The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions [3, 17, 29, 21, 31, 38].
-Milestone methods in this area are Generative Adversarial Networks (GANs) [21, 32], Variational Autoencoders (VAEs) [6, 12, 29, 38], and diffusion models [3, 19, 25, 28, 31].
+Figure 2: A comparison of generation difficulty between “toaster” and “pizza”.
+Figure 3: Comparison of easy and hard samples in a generated image.
+Figure 4: The proposed DivCon framework.
+The layout prediction stage (stage 1) is divided into numerical & spatial reasoning and bounding box planning.
+The layout-to-image generation stage (stage 2) is divided into two steps to generate objects with different difficulty levels separately.
+Related Work
+In this section, we review three main categories of text-to-image generation methods: text-to-image generation (T2I) models, layout-conditioned T2I models, and text-to-layout-to-image generation models.
+2.1 Text-to-Image (T2I) Generation Models
+The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions (Balaji et al.
+2022; Gafni et al.
+2022; Ramesh et al.
+2021; Kang et al.
+2023; Saharia et al.
+2022; Yu et al.
+2022).
+Milestone methods in this area are Generative Adversarial Networks (GANs) (Kang et al.
+2023; Sauer et al.
+2023), Variational Autoencoders (VAEs) (Chang et al.
+2023; Ding et al.
+2022; Ramesh et al.
+2021; Yu et al.
+2022), and diffusion models (Balaji et al.
+2022; Ho, Jain, and Abbeel 2020; Nichol et al.
+2022; Ramesh et al.
+2022; Saharia et al.
+2022).
 GANs often struggle with training stability and mode collapse, whereas VAEs may sometimes produce overly smooth or blurry outputs due to their reconstruction loss.
 Recently, diffusion models are highly valued for their robustness and ability to generate detailed and realistic images.
-Particularly, Stable Diffusion[30], which utilizes open vocabulary encoder[27] for encoding text to embeddings as conditional inputs, has demonstrated a remarkable capability for open-world image synthesis, pushing the boundaries of generative models.
+Particularly, Stable Diffusion (Rombach et al.
+2022), which utilizes text embeddings extracted by CLIP (Radford et al.
+2021) as conditions, has demonstrated remarkable performance for open-world image synthesis.
 Despite promising results, a text prompt is usually difficult to provide fine-grained control of the generated images.
-Figure 3: Three main categories of text-to-image generation methods.
-2.0.2 Layout-conditioned T2I Models
-To improve the controlability of T2I models, prompts like layouts as employed as additional conditions for T2I models.
+2.2 Layout-conditioned T2I Models
+To improve the controlability of T2I models, prompts like layouts are employed as additional conditions for T2I models.
 Existing layout-conditioned generation models can be categorized into three branches.
-One branch of methods trains external networks on layout-images pairs [22, 39, 25, 1, 37].
-For instance, GLIGEN incorporates a trainable gated self-attention layer that transforms bounding box information into conditional input for Stable Diffusion.
-Another branch of methods, including attention-refocusing [26], layout-predictor [34], direct-diffusion[24], layout-guidance[9] and BoxDiff [36], directly manipulate attention maps to optimize the sampling process.
-These methods generate images with multiple objects through a single forward pass of the diffusion model, resulting in limited fidelity for objects with complicated appearances or spatial relationships.
-To remedy this, the last branch of methods including Mixture-of-Diffusion[5], MultiDiffusion [4], and LLM-grounded Diffusion [23] perform denoising on each object (i.e., bounding box) separately before fusing the results.
-However, these methods introduce much higher computational costs.
-Moreover, the requirements of additional layout prompts hinder their practicality.
-2.0.3 Text-to-Layout-to-Image Generation Models
-Current LLMs have demonstrated strong numerical and spatial reasoning abilities in visual planning and layout generation.
+One branch of methods train external networks on layout-images pairs (Li et al.
+2023b; Zhang, Rao, and Agrawala 2023; Nichol et al.
+2022; Avrahami et al.
+2023; Yang et al.
+2022).
+For instance, GLIGEN incorporates a trainable gated self-attention layer that transforms bounding boxes into conditional input for Stable Diffusion.
+Another branch of methods, including attention-refocusing (Phung, Ge, and Huang 2024), layout-predictor (Wu et al.
+2023), direct-diffusion(Ma et al.
+2023), layout-guidance(Chen, Laina, and Vedaldi 2023b) and BoxDiff (Xie et al.
+2023), directly manipulate attention maps to optimize the sampling process of the diffusion model.
+Since these methods generate all objects through a single forward pass of the diffusion model, objects with spatial relationships usually suffer limited quality.
+To remedy this, the last branch of methods (including Mixture-of-Diffusion(Barbero Jiménez 2023), MultiDiffusion (Bar-Tal et al.
+2023), and LLM-grounded Diffusion (Lian et al.
+2023)) first synthesize each object (i.e., bounding box) separately and then compose them to produce the final result.
+Nevertheless, the computational cost of these methods increases with the number of objects such that the computational overhead is considerable for a text prompt containing plenty of objects.
+In addition, the relationship between objects is disrupted and cannot be well maintained in the generated images.
+2.3 Text-to-Layout-to-Image Generation Models
+Current LLMs have demonstrated strong numerical and spatial reasoning capabilities in visual planning and layout generation.
 Inspired by this, several efforts are made to integrate LLMs to generate layouts from texts, and then utilize layout-conditioned diffusion models for image generation.
-Attention-refocusing [26] and LLM-grounded Diffusion [23] explore in-context learning of GPT-4 for layout prediction.
-LayoutGPT [15] searches similar samples from their database as input examples for the LLMs.
-LLM-layout-generator [10] leverages Chain-of-Thought (CoT) prompting [33] to improve the layout prediction accuracy.
+Attention-refocusing (Phung, Ge, and Huang 2024) and LLM-grounded Diffusion (Lian et al.
+2023) explore in-context learning of GPT-4 for layout prediction.
+LayoutGPT (Feng et al.
+2023b) retrieves similar samples from a database as input examples for LLMs.
+LLM-layout-generator (Chen et al.
+2023) leverages Chain-of-Thought (CoT) prompting (Wei et al.
+2022) to improve the layout prediction accuracy.
 However, LLMs are expected to simultaneously perform numerical & spatial reasoning and visual planning of bounding boxes in these methods.
-Contrary to these methods, we propose a divide-and-conquer approach to conduct reasoning and visual planning separately, which improves the accuracy of layout predictions and substantially advances the prompt fidelity of generated images.
-In addition, during layout-to-image generation, the objects are synthesized in an iterative manner according to their difficulties, which achieves not only superior fidelity but also significant computational cost savings.
-Figure 4: The proposed DivCon framework.
-The layout prediction stage is divided into numerical & spatial reasoning and bounding box planning.
-The layout-to-image generation is divided into two iterations to generate objects with different difficulty levels separately.
-3 Method
-3.1 Overview
+As a result, inferior accuracy is produced especially for complicated text prompts.
+Contrary to these methods, we propose a divide-and-conquer approach to conduct reasoning and layout prediction separately, which improves the accuracy of layout predictions and substantially advances the prompt fidelity of generated images.
+In addition, during layout-to-image generation, the objects are synthesized progressively according to their difficulties, which achieves not only superior fidelity but also significant computational cost savings.
+Methods Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc. Spatial Layout Image Acc. Acc.
+Stable Diffusion 2.1 - - - - 73.97 53.93 62.38 16.20 - 7.96
+Attend-and-Excite - - - - 76.57 55.94 64.65 18.37 - 9.23
+Divide-and-Bind - - - - 77.95 57.50 66.18 20.15 - 13.15
+Our layout + Layout Guidance - - - - 71.09 45.98 55.84 16.58 - 29.85
+Our layout + GLIGEN - - - - 75.34 62.14 69.01 23.72 - 48.41
+Attention-Refocusing 88.62 91.91 90.44 77.05 77.43 61.78 68.72 25.60 89.57 48.29
+DivCon (Ours) 94.17 89.53 91.80 80.73 78.65 63.41 70.22 29.97 91.87 54.21
+Table 1: Comparison of our method with baselines on HRS benchmark. Text in bold shows the best results of each metric.
+Method
 As illustrated in Fig.
 4, our framework comprises two stages.
-Given a text prompt, LLMs first predict layouts from it.
-Then, the resultant layout is employed as an additional condition to achieve text-to-image generation.
+Given a text prompt, LLMs is first adopted to parse it and predict a layout.
+Then, the resultant layout is fed to the diffusion model as an additional condition with the text prompt to achieve text-to-image generation.
 More specifically, each stage is further divided into two steps to handle simple subtasks.
 3.1 Text-to-Layout Prediction
-Intuitively, to predict the layout from a text prompt, the number & spatial relations of the objects are first required to be reasoned, and then bounding boxes are generated to describe the coordinates & sizes of the objects.
+Intuitively, to predict the layout from a text prompt, the number & spatial relations of the objects are first required to be reasoned, and then bounding boxes are generated to describe the coordinates and sizes of the objects.
 However, previous methods commonly conduct these two steps simultaneously without considering their inherent correlation, thereby hindering them from understanding complicated text prompts well.
-To remedy this, we introduce the idea of divide-and-conquer to improve the accuracy of predicted layouts.
-Specifically, leveraging the state-of-the-art large language model GPT-4 [], numerical and spatial reasoning is first performed, and then bounding boxes are predicted.
-3.2.1 Numerical and Spatial Reasoning
-Given a text prompt $C$ describing an image, an LLM is adopted to predict a set of strings $O=\{s_{j}|j=1,2,...,n\}$.
-If the prompt primarily focused on counting relationships among objects, each string $s_{j}$ includes both the name and quantity information of each object category $j$.
-If the prompt mainly describes spatial relationships, each string $s_{j}$ will contain the object name and its spatial position in the image.
-This method allows LLMs to conclude complex numerical and spatial relationships.
-3.2.2 Bounding Box Prediction
-After numerical and spatial reasoning, both the meta prompt $C$ and the prediction result $O$ are fed to the LLM again to predict a list of name-box pairs $P=\{p_{i}|i=1,2,...,n\}$.
+To remedy this, we employ the divide-and-conquer strategy to improve the accuracy of predicted layouts.
+Specifically, the text prompt is first fed to an LLM (e.g., GPT-4 (OpenAI 2024)) to perform numerical and spatial reasoning to produce more distinct instructions of objects, as illustrated in Fig.
+4(a).
+Then, these instructions are adopted as additional conditions to predict more accurate bounding boxes.
+(1) Numerical and Spatial Reasoning.
+Given a text prompt $C$ describing an image, an LLM is adopted to predict a set of strings $O=\{s_{j}|j=1,2,...,n\}$ to describe the objects in the prompt.
+If the prompt primarily focuses on counting relationships among objects, each string $s_{j}$ includes both the name and quantity information of each object category $j$.
+If the prompt mainly describes spatial relationships, each string $s_{j}$ contains the object name and its spatial position in the image.
+After numerical and spatial reasoning, the numerical information and spatial relationships can be explicitly parsed from the complicated text prompt as intermedium conditions.
+(2) Bounding Box Prediction.
+After numerical and spatial reasoning, both the meta prompt $C$ and intermedium conditions $O$ are fed to the LLM again to predict a list of name-box pairs $P=\{p_{i}|i=1,2,...,n\}$.
 Each pair $p_{i}=\{o_{i}:(l_{i},t_{i},r_{i},b_{i})|i=1,2,...,n\}$ consists of the object name $o_{i}$ and its bounding box’s coordinates $(l_{i},t_{i},r_{i},b_{i})$.
-The in-context exemplars include a process of correcting the features of generated bounding boxes to ensure the quality of the output layouts.
+Meanwhile, an in-context exemplars including a process of correcting the features of generated bounding boxes is provided to the LLM to improve the quality of the predicted layouts.
 3.2 Layout-to-Image Generation
 After layout prediction, the resultant layout is employed as an additional condition to generate the image.
-As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy to hard in an iterative manner, as shown in Fig.
-4.
-First, the layout is fed to the diffusion model to synthesize an image.
+As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy ones to hard ones in an progressive manner, as shown in Fig.
+4(b).
+First, the layout is fed to the diffusion model to synthesize all objects.
 Then, the consistency between the synthetic objects and the text prompt is calculated.
-Afterward, objects with low fidelity are highlighted and passed to the diffusion model for a second time, encouraging the diffusion model to focus on these difficult objects to achieve higher quality.
-3.3.1 First-Round Generation
-We employ a layout-to-image diffusion model with sampling optimization [26] as the generative model to generate an image conditioned on the text prompt and the predicted layout prompt.
-During the denoising process, random noise $z_{T}$ is fed to the diffusion model to gradually denoise it to an image $z_{0}$ after $T$ iterations.
-3.3.2 Consistency Evaluation
-With the generated image $x$, each object is cropped using the bounding boxes in the layout, which are denoted as $x_{i}\in\mathbb{R}^{4\times h\times w}$, where $h\times w$ represents the height and width of the bounding box.
-Then, the CLIP similarity $S_{x_{i},o_{i}}\in(-1,1)$ is employed to evaluate the consistency between the cropped patch and the object name $o_{i}$ in the text prompt.
+Objects with low consistency are considered as low-fidelity ones and fed to the diffusion model for a second time while maintaining the results of objects with high consistency.
+With low-fidelity objects being highlighted, the diffusion model is encouraged to focus on these difficult objects in the second forward pass to achieve higher quality.
+(1) First-Round Generation.
+We employ a layout-to-image diffusion model with sampling optimization (Phung, Ge, and Huang 2024) as the generative model to generate an image conditioned on the text prompt and the predicted layout.
+During the denoising process, random noise $z_{T}$ is fed to the diffusion model to gradually denoise it to an image $z_{0}$ after $T$ time steps.
+(2) Consistency Evaluation.
+In the generated image $z_{0}$, each object is cropped using the bounding boxes in the layout and denoted as $x_{i}\in\mathbb{R}^{4\times h\times w}$, where $h\times w$ represents the height and width of the bounding box.
+Then, the CLIP similarity $S_{x_{i},o_{i}}\in(-1,1)$ is employed to evaluate the consistency between the cropped object and its name $o_{i}$ in the text prompt.
 The higher the similarity is, the higher the fidelity of the corresponding object is.
-Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are "good enough”.
-In contrast, other objects with low similarities are hard samples that are more difficult to reconstruct (e.g.
-"sheep" and "horse" in Fig.
-5).
-Next, a binary mask $M$ is calculated with bounding boxes of easy objects being set to 1 and the rest regions being set to 0.
-Figure 5: Comparison of easy and hard samples in a generated image.
-3.3.3 Second-Round Refinement
-After consistency evaluation, different objects are of various levels of fidelity.
-Our objective is to preserve the “good enough” objects while reproducing the remaining ones to ease the difficulty of diffusion model.
-To this end, we retrieve the latent representation during the denoising process of the first-round generation.
-Specifically, $z_{0.7T}$ and $z_{0.3T}$ are first retrieved.
-Then, the binary mask $M$ is adopted to crop high-fidelity regions in $z_{0.7T}$ and $z_{0.3T}$.
-Next, the starting noise $z^{\prime}_{T}$ for the second-round diffusion is obtained as:
-$z^{\prime}_{T}=z_{0.7T}\circ M+\mathcal{N}(\mu,\sigma^{2})\circ(1-M)$ (1)
-where $\circ$ is element-wise multiplication and $\mathcal{N}(\mu,\sigma^{2})$ is random noise.
-The second-round denoising process consists of three stages.
-First, the input noise $z^{\prime}_{T}$ is denoised for $0.3T$ iterations to obtain $z^{\prime}_{0.3T}$.
-Second, the high-fidelity regions in $z_{0.7T}$ are cropped and used to replace the corresponding regions in $z^{\prime}_{0.3T}$ and fed to the diffusion model for anthoer $0.4T$ denoising iterations.
-Note that, the text prompt and the full layout prompt are employed as conditions in these two stages.
-Moreover, the regions marked by $M$ is frozen to preserve these high-fidelity regions.
-During optimization, Cross-Attention Refocusing (CAR) and Self-Attention Refocusing (SAR) losses [26] are employed to update only regions of low fidelity, encouraging the model to focus on these hadr samples.
-Finally, all regions are optimized for the last $0.3T$ iterations to smoothly blend the high-fidelity regions with other regions.
-4 Experiments
+Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are ”good enough”.
+In contrast, other objects with low similarities are considered as hard samples that are more difficult to synthesize (e.g.
+”sheep” and ”horse” in Fig.
+3).
+Next, a binary mask $M$ is produced with bounding boxes of easy objects set to 1 and the rest regions set to 0.
+(3) Second-Round Refinement.
+The objective of second-round refinement is to preserve the “good enough” objects while reproducing the remaining ones to ease the difficulty of the diffusion model.
+To this end, a straightforward approach is to retrieve the latent representation of the objects in the first-round denoising process.
+Nevertheless, directly freezing the object regions during the denoising process produces inharmonious artifacts around the boundaries.
+To address this issue, we retrieve the latent representation $z_{T}$ of all time steps during the denoising process of the first-round generation.
+Then, the binary mask $M$ is adopted to crop high-fidelity regions in the latent representation.
+At each time step $T$ of the second-round forward pass, these cropped representations are employed to replace the corresponding regions in the noises:
+$z^{\prime}_{T}=z_{T}\circ M+z^{\prime}_{T}\circ(1-M)$ (1)
+where $\circ$ is element-wise multiplication and $z^{\prime}_{T}$ is the noise at $T$ step of the second-round forward pass.
+With the latent representation of high-fidelity objects (masked by $M$) being retrieved, these objects can be well preserved to encourage the model to focus on reconstructing the remaining ones.
+Meanwhile, the bounding boxes of “bad samples” $P^{\prime}$ are passed to the diffusion model as the layout conditions, enabling the model to concentrate more on refining $P^{\prime}$ to further improve the overall quality of the generated image.
+Methods Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc. Spatial Layout Image Acc. Acc.
+Stable Diffusion 2.1 - - - - 78.94 75.42 77.41 33.99 - 22.97
+Attend-and-Excite - - - - 79.21 78.62 78.91 39.51 - 29.61
+Divide-and-Bind - - - - 84.99 79.13 81.96 32.35 - 24.00
+Our layout + Layout Guidance - - - - 81.97 75.41 78.55 43.13 - 52.88
+Our layout + GLIGEN - - - - 85.13 84.97 85.05 52.21 - 59.72
+Attention-Refocusing 98.51 99.11 98.81 94.31 84.61 85.64 85.12 56.10 95.05 69.28
+DivCon (Ours) 98.47 98.35 98.41 94.39 85.41 87.06 86.22 57.35 96.23 73.15
+Table 2: Comparison of our method with baselines on NSR-1K benchmark. Text in bold shows the best results of each metric.
+Experiments
 In this section, we provide an evaluation of our method and compare it with state-of-the-art T2I models.
-Ablation experiments are conducted to demonstrate the effectiveness of our divide-and-conquer approach in layout prediction and layout-to-image generation stages.
+Ablation experiments are then conducted to demonstrate the effectiveness of our divide-and-conquer approach in layout prediction and layout-to-image generation stages.
 4.1 Experiment Setup
-4.1.1 Benchmarks
-To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS[2] and NSR-1K[15].
+(1) Benchmarks and Baselines
+To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS (Bakr et al.
+2023) and NSR-1K (Feng et al.
+2023b).
 The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships.
 The NSR-1K dataset consists of 762 prompts for numerical relationships and 283 prompts for spatial relationships, all labeled with names, counts, bounding boxes, and ground truth images.
-We also utilize NSR-1K for quality evaluation.
 Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations.
-This elevates HRS as a more sophisticated and challenging benchmark.
-4.1.2 Evaluation Metrics
-We evaluate the images produced by our methods in terms of both quality and fidelity.
-(1) Quality To assess the quality of generated images, we use FID as the metric[18].
-Lower FID represents higher similarity between generated images and ground truths images.
-We report FID in NSR-1K [15] as each prompt has a corresponding ground truth image for FID calculation.
-(2) Fidelity To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated.
-Particularly, YOLOv8 [20] is employed to obtain the bounding boxes of the objects in the generated images.
-In counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy [13, 16].
-In spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills[11] and LayoutGPT [15].
+For these benchmarks, Stable Diffusion (Rombach et al.
+2022), Attend-and-Excite (Chefer et al.
+2023), Divide-and-Bind (Li et al.
+2023a), GLIGEN (Li et al.
+2023b), Layout-Guidance (Chen, Laina, and Vedaldi 2023b), and Attention-Refocusing (Phung, Ge, and Huang 2024) are included for comparison.
+In addition, we also include the dataset in (Lian et al.
+2023) for evaluation, which includes four tasks: negation, generative numeracy, attribute binding, and spatial reasoning.
+As our method focuses on numerical and spatial reasoning, we select 100 prompts from these two tasks to construct a benchmark termed LLM-Grounded.
+(2) Evaluation Metrics
+Quality. To assess the quality of generated images, we use FID as the metric (Heusel et al. 2017). Lower FID represents higher similarity between generated images and ground truths.
+Fidelity.
+To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated.
+Particularly, YOLOv8 (Jocher, Chaurasia, and Qiu 2023) is employed to obtain the bounding boxes of the objects in the generated images.
+For counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy (El-Nouby et al.
+2019; Fu et al.
+2020).
+For spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills (Cho, Zala, and Bansal 2022) and LayoutGPT (Feng et al.
+2023b).
 Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship.
 Otherwise, they are denoted as a top-bottom relationship.
-4.1.3 Baselines
-Stable Diffusion [30], Attend-and-Excite[7], and Attention-Refocusing [26] are employed for comparison.
-Stable Diffusion and Attend-and-Excite directly generate images from the text prompt.
-Meanwhile, Attention-Refocusing employs layout as an intermedium for image generation.
-Table 1: Comparison of our method with baselines on HRS benchmark. Text in bold shows the best results of each metric.
-Methods
-Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc.
-Spatial
-Layout Image
-Acc.
-Acc.
-Stable Diffusion 2.1[30]
-- - - - 73.97 53.93 62.38 16.20
-7.96
-Attend-and-Excite[7]
-- - - - 76.57 55.94 64.65 18.37
-9.23
-Attention-Refocusing[26]
-88.62 91.91 90.44 77.05 77.43 61.78 68.72 25.60
-89.57
-48.29
-DivCon (Ours)
-94.17 89.53 91.80 80.73 79.13 63.53 70.48 29.62
-91.87
-53.87
-Table 2: Comparison of our method with baselines on NSR-1K benchmark. Text in bold shows the best results of each metric.
-Methods
-Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc.
-Spatial
-Layout Image
-Acc.
-Acc.
-Stable Diffusion 2.1[30]
-- - - - 78.94 75.42 77.41 33.99
-22.97
-Attend-and-Excite[7]
-- - - - 79.21 78.62 78.91 39.51
-29.61
-Attention-Refocusing[26]
-98.51 99.11 98.81 94.31 84.61 85.64 85.12 56.10
-95.05
-69.28
-DivCon (Ours)
-98.47 98.35 98.41 94.39 85.31 87.54 86.41 55.80
-96.11
-72.95
-Figure 6: Qualitative comparison of numerical reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing, and our DivCon.
-Figure 7: Qualitative comparison of spatial reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing and our DivCon.
-Table 3: FID results achieved by different methods in NSR-1K.
-Methods
-FID ($\downarrow$)
-Stable Diffusion 2.1[30]
-20.77
-Attention-Refocusing[26]
-21.05
-DivCon (Ours)
-20.89
+Efficiency. To validate the efficiency and scalability of our method, inference runtime is evaluated on an AWS g5.xlarge GPU instance.
+Figure 5: Qualitative comparison of numerical & spatial reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing and our DivCon.
 4.2 Evaluation Results
-4.2.1 Quantitative Results
-We present the results on the HRS benchmark [2] in Tab.
+(1) Quantitative Results
+The HRS dataset.
+We present the results on the HRS benchmark in Tab.
 1.
-As we can see, DivCon outperforms previous T2I models by over 10% in numerical accuracy and over 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance.
-In addition, DivCon surpasses Attention Refocusing [26] on most metrics.
+As we can see, our DivCon outperforms previous T2I models by over 10% in numerical accuracy and about 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance.
+In addition, DivCon surpasses Attention Refocusing (Phung, Ge, and Huang 2024) on most metrics.
 Specifically, our layout prediction improves numerical accuracy by 3% and spatial accuracy by 2%, respectively.
@@
 By dividing the T2I task into simple subtasks, our DivCon can better understand the text prompt to achieve higher fidelity.
+The NSR-1K dataset.
 The results on the NSR-1K benchmark are presented in Tab.
 2.
-It can be observed that DivCon outperforms other T2I models by over 15% in numerical accuracy and over 45% in spatial accuracy.
-As compared to Attention Refocusing [26], our DivCon produces competitive results in terms of layout accuracy.
-Meanwhile, DivCon achieves approximately 2% superiority in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy.
+It can be observed that DivCon outperforms other T2I models by over 20% in numerical accuracy and about 45% in spatial accuracy.
+As compared to Attention Refocusing (Phung, Ge, and Huang 2024), our DivCon produces competitive results in terms of layout accuracy.
+Meanwhile, DivCon achieves approximately 2% gains in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy.
 Quantitative results in terms of image quality are shown in Tab.
 3.
-As we can see, our DivCon produces compatible FID scores as compared to stable diffusion and outperforms attention-refocusing with notable margins.
-This demonstrates that our approach is able to preserve image quality while enhancing prompt fidelity.
-4.2.2 Qualitative Results
-Fig. 6 and Fig. 7 illustrate the qualitative comparison of DivCon and baselines. In all cases, our DivCon can accurately predict the object quantities and the spatial relations from the text prompts.
-(1) Large Quantity of Objects.
+As we can see, our DivCon produces comparable FID scores to Stable Diffusion and outperforms attention-refocusing with notable margins.
+This further demonstrates that our approach is able to produce images with better visual quality while maintaining high prompt fidelity.
+Methods FID ($\downarrow$)
+Stable Diffusion 2.1 20.77
+Attention-Refocusing 21.05
+DivCon (Ours) 21.01
+Table 3: FID results achieved on NSR-1K.
+LLM-Grounded benchmark.
+Table 4 presents the accuracy achieved by our method and LLM-grounded Diffusion.
+It is clear that our method achieves comparable performance to LLM-Grounded Diffusion on the numerical reasoning while outperforming it by 3% on the spatial reasoning task.
+This also validates the superiority of our method.
+Methods Numerical Spatial
+LLM-Grounded (GPT4) 84% 82%
+DivCon (Ours) 84% 87%
+Table 4: Comparison with LLM-Grounded
+Runtime comparisons are presented in Tab.
+5.
+Among techniques that require multiple forward passes to synthesize different regions separately, our DivCon produces notable efficiency gains.
+Particularly, LLM-Grounded takes 134.98s to generate an image, while our method achieves a 1.87×1.87$\times$1.87 × speedup.
+Compared with methods with a single forward pass, our method (1 Iter) produces competitive results with comparable efficiency.
+Methods #Forward Pass Runtime
+Stable Diffusion 1 15.81 s
+GLIGEN 1 33.75 s
+Attention-Refocusing 1 35.71 s
+DivCon-1 iter (Ours) 1 37.35 s
+Mixture-of-Diffusion 3 126.41 s
+LLM-Grounded 3 134.98 s
+Multi-Diffusion 3 121.68 s
+DivCon-2 iter (Ours) 2 72.01 s
+Table 5: Comparison of Inference Runtime.
+(2) Qualitative Results
+Figure 5 illustrates the qualitative comparison of DivCon and baselines. In all cases, our DivCon can accurately predict the object quantities and the spatial relations from the text prompts.
+Large Quantity of Objects.
 Given a prompt containing more than one object category (Fig.
-6), Stable Diffusion [30] and Attend-and-Excite [7] often generate an insufficient quantity of objects, while Attention-Refocusing [26] sometimes omits one category of objects or generates an excessive number of objects.
+5), Stable Diffusion (Rombach et al.
+2022) and Attend-and-Excite (Chefer et al.
+2023) often generate an insufficient quantity of objects, while Attention-Refocusing (Phung, Ge, and Huang 2024) sometimes omits one category of objects or generates an excessive number of objects.
 For instance, the prompt in Fig.
-6(b) contains five spoons and four oranges.
+5(b) contains five spoons and four oranges.
 However, Stable Diffusion and Attend-and-Excite cannot synthesize sufficient spoons or oranges.
@@
 As compared to these methods, our DivCon achieves superior numerical accuracy.
-(2) Complicated Spatial Relation.
-Given a prompt containing complex spatial relationships among more than two object categories (Fig.
-7, previous T2I models struggle to generate correct spatial relations.
-Meanwhile, Attention-Refocusing tends to produce distortion or blending artifacts around the boundaries of objects.
-In Fig.
-7 (b), Stable Diffusion and Attend-and-Excite miss certain object categories, while Attention-Refocusing merges the color of the banana with the person.
-In contrast, our DivCon inference accurate object categories and their spatial relationships, exhibiting superior performance in these challenging cases.
-Figure 8: Comparison of layouts generated from Attention-Refocusing and our DivCon.
-Figure 9: Ablation study of our image generation.
-Figure 10: Comparison of transmit one or two latent representations.
+Complicated Spatial Relation.
+Given a prompt containing complex spatial relationships among more than two object categories, previous T2I models struggle to generate correct spatial relations.
+For example, Stable Diffusion and Attend-and-Excite miss certain object categories (Fig.
+5(d)).
+Meanwhile, Attention-Refocusing tends to produce distortion or blending artifacts around the boundaries of objects (Fig.
+5(c)(d)) or merges the color of the banana with the person (Fig.
+5(d)).
+In contrast, our DivCon produces accurate object categories and spatial relationships, exhibiting superior performance in these challenging cases.
+User Study.
+We conduct a user study to evaluate the images generated by different methods.
+To be specific, we recruit 17 people with different backgrounds and asked them to rank the images generated by different methods according to their fidelity to the original prompt.
+Lower ranks indicates better alignment.
+A total of 20 examples are randomly selected from the NSR-1K benchmark.
+As shown in Fig.
+6, our method achieves the best performance in terms ofaverage rank, which demonstrates the superior perceptual quality of our results.
+Figure 6: User Study: Comparison of image quality.
 4.3 Ablation Study
@@
 Results are presented in Tab.
-4 and Attention-Refocusing [26] is used as the baseline.
-4.3.1 Divide-and-Conquer in Layout Prediction
+6 and Attention-Refocusing (Phung, Ge, and Huang 2024) is used as the baseline.
+Divide-and-Conquer in Layout Prediction.
 To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step.
@@
 We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig.
-8.
+7.
 Compared to Attention-Refocusing, layouts generated by our DivCon not only exhibit higher reasoning accuracy (Fig.
-8 (a), (b)) but also are more organized with less overlap (Fig.
-8 (c), (d)).
-4.3.2 Divide-and-Conquer in Image Generation
+7 (a) but also are more organized with less overlap (Fig.
+7 (b)).
+Layout w/ DivCon Image w/ DivCon Numerical Precision Recall F1 Acc. Spatial Acc.
+(1) - - 77.43 61.78 68.72 25.60 48.29
+(2) - 78.26 62.61 69.57 29.46 52.74
+(3) - 78.18 62.36 69.38 27.83 53.17
+(4) 79.13 63.53 70.48 29.62 53.87
+Table 6: Ablation study of our DiVCon on HRS benchmark.
+“Layout w/ DivCon”: Layout prediction with divide-and-conquer strategy.
+“Image w/ DivCon”: Layout-to-Image generation with divide-and-conquer strategy.
+Figure 7: Comparison of the predicted layouts.
+Figure 8: Ablation study of our image generation.
+Divide-and-Conquer in Image Generation.
 To study the effectiveness of our divide-and-conquer approach in the image generation stage, we develop a variant to synthesize the image in a single forward pass.
@@
 We further compare the results produced by our method using prompts containing multiple objects in Fig.
-9.
-As we can see, the divide-and-conquer approach facilitates our DivCon to better reconstruct the objects, especially for small-sized ones, while avoiding generating objects outside the designated bounding boxes.
-For example, in Fig.
-9(a), DivCon accurately reconstructs the small objects.
-In Fig.
-9(d), DivCon successfully avoids generating the dog outside boxes.
-Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy.
-4.3.3 Fractiles in Image Generation
-During image generation of our DivCon, two fractiles (i.e., $z_{0.3T}$ and $z_{0.7T}$) are used to preserve “good enough” results in the first-round generation.
-We conduct ablation experiments to study these two fractiles.
-The second fractile $z_{0.7T}$ marks the end of optimization and is non-adjustable.
-We observe that transmitting only $z_{0.7T}$ to the next phase results in instability in image generation instability, specifically manifesting as noticeable boundary demarcations of bounding boxes in the images (Fig.
-10).
-We further compare the performance achieved by our model with different values for the first fractile.
-Experiment results are presented in Tab.
-5.
-It can be observed that the model suffers a notable performance drop if only the second fractile is employed.
-With an additional fractile, our DivCon yields a notable improvement.
-Additionally, the selection of the first fractile’s value exhibits a marginal impact on the efficacy of the model, with the 30% threshold demonstrating optimal performance.
-Table 4: Ablation study of our method’s image generation performance on HRS benchmark.
-"Layout w/DivCon": Layout prediction from LLMs with divide-and-conquer prompting.
-"Image w/DivCon": Layout-to-Image generation with divide-and-conquer strategy.
-Layout w/DivCon Image w/DivCon
-Numerical
-Precision
-Recall
-F1
-Acc.
-Spatial
-Acc.
-77.43
-61.78
-68.72
-25.60
-48.29
-78.26
-62.61
-69.57
-29.46
-52.74
-78.18
-62.36
-69.38
-27.83
-53.17
-79.13
-63.53
-70.48
-29.62
-53.87
-Table 5: Ablation study of two fractiles used to preserve results in the first-round generation.
-First Fractile Second Fractile
-Numerical
-Precision
-Recall
-F1
-Acc.
-Spatial
-Acc.
-76.82
-60.69
-67.80
-24.31
-48.53
-$0.1T$
-78.44
-61.61
-69.01
-29.12
-52.30
-$0.3T$
-79.13 63.53 70.48 29.62
-53.87
-$0.5T$
-78.13
-63.25
-69.90
-28.71
-51.46
-5 Limitations and Discussions
+8.
+As we can see, the divide-and-conquer approach facilitates our method to better reconstruct the objects.
+For example, DivCon accurately reconstructs the small objects (Fig.
+8(a)) and avoids generating the dog outside boxes (Fig.
+8(b)).
+Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy (Tab.
+6).
+4.4 Limitations and Discussions
 As illustrated in Fig.
-11, our method still struggles with generating objects from certain pattern layouts.
-When the text prompt describes objects overlapping in space, the diffusion model may fail to generate all objects accurately.
+9, when the text prompt describes objects with extreme overlap, the diffusion model may fail to generate all objects accurately.
 This limitation stems from the inherent capacity of layout-conditioned image generation models to handle overlapping boxes.
-Despite these challenges, our approach still demonstrates a reduced frequency of overlapping in layout prediction compared to the base model.
-Figure 11: Comparison of transmit one or two latent representations.
-6 Conclusion
-In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into simple subtasks to achieve higher image generation fidelity.
+Despite these challenges, our approach still demonstrates an improved quality in layout prediction as compared to the base model.
+Figure 9: An illustration of failure case.
+Conclusion
+In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into multiple subtasks to achieve higher image generation quality.
 Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, respectively.
-Then, layout-to-image generation is achieved via an iterative manner to synthesize objects in an easy-to-hard fashion.
+Then, layout-to-image generation is achieved via an progressive manner to synthesize objects in an easy-to-hard fashion.
 Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.

### v3

@@
-DivCon: Divide and Conquer for Progressive Text-to-Image Generation
+DivCon: Divide and Conquer for Complex Numerical and Spatial Reasoning in Text-to-Image Generation
 Abstract
@@
 To further improve T2I models’ capability in numerical and spatial reasoning, layout is employed as an intermedium to bridge large language models and layout-based diffusion models.
-However, these methods still struggle with generating images from prompts with multiple objects and complicated spatial relationships.
-To tackle this challenge, we introduce a divide-and-conquer approach which decouples the generation task into multiple subtasks.
-First, the layout prediction stage is divided into numerical & spatial reasoning and bounding box prediction to better parse the text prompt to obtain accurate composition information.
+However, these methods often rely on closed-source, large-scale LLMs for layout prediction, limiting accessibility and scalability.
+They also struggle with generating images from prompts with multiple objects and complicated spatial relationships.
+To tackle these challenges, we introduce a divide-and-conquer approach which decouples the generation task into multiple subtasks.
+First, the layout prediction stage is divided into numerical & spatial reasoning and bounding box visual planning, enabling even lightweight LLMs to achieve layout accuracy comparable to large-scale models.
 Second, the layout-to-image generation stage is divided into two steps to synthesize objects from easy ones to difficult ones.
@@
 In addition, visual results and user study demonstrate that our approach significantly improves the perceptual quality, especially when generating multiple objects from complex textural prompts.
-Codehttps://github.com/DivCon-gen/DivCon
-Introduction
-Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects (Rombach et al.
-2022; Chefer et al.
-2023; Balaji et al.
-2022; Kang et al.
-2023; Ramesh et al.
-2021; Saharia et al.
-2022).
-However, these methods still suffer limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships (Bakr et al.
-2023; Chefer et al.
-2023; Feng et al.
-2023a; Chen, Laina, and Vedaldi 2023a; Xiao et al.
-2023).
-Recently, combining large language models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ capability in numerical and spatial reasoning (Feng et al.
-2023b; Phung, Ge, and Huang 2024).
+[…]
+0783
+1 Introduction
+Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects [34, 7, 3, 22, 32, 35].
+However, these methods still suffer limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships [2, 7, 13, 8, 40].
+Recently, combining large language models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ capability in numerical and spatial reasoning [14, 29].
 These methods leverage the strong reasoning and visual planning ability of LLMs to predict objects’ layouts from the text prompt, which provides additional informative conditions for diffusion models to produce higher-quality images.
-Despite recent significant advancements, existing methods still suffer limited accuracy when the text prompt contains multiple objects and complicated spatial relationships (Bakr et al.
-2023; Phung, Ge, and Huang 2024).
-On the one hand, LLMs encounter great challenges when conducting simultaneous reasoning and visual planning from complicated text prompts (e.g.
-“five mice are scurrying around a wooden table, each nibbling on one of the five ripe apples that are laid out on top”).
-On the other hand, since the difficulty of synthesizing objects with different characteristics varies a lot, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and details.
-However, previous methods commonly conduct reasoning and visual planning together while generating all objects simultaneously without taking their differences into consideration.
-As a result, objects with higher difficulties cannot be well synthesized in the images (e.g.
-toaster and pizza in Fig.
-2).
+Despite recent significant advancements, existing methods still suffer key limitations when dealing with complex prompts involving multiple objects and complex spatial constraints [2, 29].
+Two major factors hinder their practical effectiveness.
+First, most methods rely on large-scale LLMs such as GPT-4 [28] to predict layouts[14, 29, 25, 30], which, while effective, requires substantial computational resources and limited accessibility.
+In contrast, whether lightweight, open-source, pretrained language models can perform accurate layout reasoning remains largely unexplored.
+Second, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and detail complexity.
+However, most existing methods synthesize all objects simultaneously, without accounting for the varying difficulty of generating different object types.
+As a result, more complex objects are often poorly synthesized or generated with incorrect counts and spatial positions (e.g., Fig.
+2), leading to visible inconsistencies in the final images.
 Figure 1: Layouts and images generated by our DivCon. DivCon enhances the capability of text-to-image diffusion models to understand complex numerical and spatial relationships in the text.
-To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation stages.
-During layout prediction, the number of objects and their spatial relationships are first reasoned from the text prompt.
-Then, visual planning is conducted to predict the bounding boxes of the objects.
-By decoupling layout prediction into reasoning and visual planning, more accurate composition information can be obtained.
+To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation stages (Fig.
+1).
+In the layout prediction stage, we leverage a lightweight, open-source language model to first perform numerical and spatial reasoning from the text prompt, extracting object counts and positional relations.
+We then use these reasoning results to guide visual planning through constrained decoding.
+This decoupled structure significantly improves the numerical and spatial accuracy of the predicted layouts.
+Despite using only smaller-scale models, our approach achieves layout prediction performance comparable to methods relying on large LLMs, while requiring far fewer computational resources.
 During layout-to-image generation, the diffusion model is first executed to process the layout to generate all objects.
@@
 By diving layout-to-image generation into two steps, objects with different levels of difficulties can be synthesized progressively for higher quality.
-Extensive experiments have demonstrated that our DivCon has significantly improved the quality of generated images and achieves state-of-the-art performance on the HRS(Bakr et al.
-2023) and NSR-1K(Feng et al.
-2023b) benchmark datasets.
+Extensive experiments have demonstrated that our method DivCon has significantly improved the quality of generated images and achieves state-of-the-art performance on the HRS[2] and NSR-1K[14] benchmark datasets.
 Our main contributions can be summarized as follows:
   • We propose DivCon, a divide-and-conquer approach for layout-based text-to-image generation by dividing the complicated task into multiple subtasks.
-  • To obtain accurate composition information from the text prompt, we divide the layout prediction stage into numerical & spatial reasoning and bounding box planning.
+  • To improve layout prediction accuracy with lightweight LLMs to reduce computational cost, we divide layout prediction into numerical & spatial reasoning and bounding box planning.
   • To generate high-fidelity images from the layout, we divide layout-to-image generation into two steps to synthesize objects with different levels of difficulty.
   • Experimental results show that our approach achieves significant performance gains, surpassing current state-of-the-art models.
-Figure 2: A comparison of generation difficulty between “toaster” and “pizza”.
+Figure 2: A comparison of generation difficulty between “toaster" and “pizza".
 Figure 3: Comparison of easy and hard samples in a generated image.
@@
 The layout-to-image generation stage (stage 2) is divided into two steps to generate objects with different difficulty levels separately.
-Related Work
+2 Related Work
 In this section, we review three main categories of text-to-image generation methods: text-to-image generation (T2I) models, layout-conditioned T2I models, and text-to-layout-to-image generation models.
 2.1 Text-to-Image (T2I) Generation Models
-The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions (Balaji et al.
-2022; Gafni et al.
-2022; Ramesh et al.
-2021; Kang et al.
-2023; Saharia et al.
-2022; Yu et al.
-2022).
-Milestone methods in this area are Generative Adversarial Networks (GANs) (Kang et al.
-2023; Sauer et al.
-2023), Variational Autoencoders (VAEs) (Chang et al.
-2023; Ding et al.
-2022; Ramesh et al.
-2021; Yu et al.
-2022), and diffusion models (Balaji et al.
-2022; Ho, Jain, and Abbeel 2020; Nichol et al.
-2022; Ramesh et al.
-2022; Saharia et al.
-2022).
+The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions [3, 16, 32, 22, 35, 43].
+Milestone methods in this area are Generative Adversarial Networks (GANs) [22, 36], Variational Autoencoders (VAEs) [6, 11, 32, 43], and diffusion models [3, 18, 27, 33, 35].
 GANs often struggle with training stability and mode collapse, whereas VAEs may sometimes produce overly smooth or blurry outputs due to their reconstruction loss.
 Recently, diffusion models are highly valued for their robustness and ability to generate detailed and realistic images.
-Particularly, Stable Diffusion (Rombach et al.
-2022), which utilizes text embeddings extracted by CLIP (Radford et al.
-2021) as conditions, has demonstrated remarkable performance for open-world image synthesis.
+Particularly, Stable Diffusion [34], which utilizes text embeddings extracted by CLIP [31] as conditions, has demonstrated remarkable performance for open-world image synthesis.
 Despite promising results, a text prompt is usually difficult to provide fine-grained control of the generated images.
 2.2 Layout-conditioned T2I Models
-To improve the controlability of T2I models, prompts like layouts are employed as additional conditions for T2I models.
+To improve the controllability of T2I models, prompts like layouts are employed as additional conditions for T2I models.
 Existing layout-conditioned generation models can be categorized into three branches.
-One branch of methods train external networks on layout-images pairs (Li et al.
-2023b; Zhang, Rao, and Agrawala 2023; Nichol et al.
-2022; Avrahami et al.
-2023; Yang et al.
-2022).
+One branch of methods train external networks on layout-images pairs [24, 44, 27, 1, 42].
 For instance, GLIGEN incorporates a trainable gated self-attention layer that transforms bounding boxes into conditional input for Stable Diffusion.
-Another branch of methods, including attention-refocusing (Phung, Ge, and Huang 2024), layout-predictor (Wu et al.
-2023), direct-diffusion(Ma et al.
-2023), layout-guidance(Chen, Laina, and Vedaldi 2023b) and BoxDiff (Xie et al.
-2023), directly manipulate attention maps to optimize the sampling process of the diffusion model.
+Another branch of methods, including attention-refocusing [29], layout-predictor [39], direct-diffusion[26], layout-guidance[8] and BoxDiff [41], directly manipulate attention maps to optimize the sampling process of the diffusion model.
 Since these methods generate all objects through a single forward pass of the diffusion model, objects with spatial relationships usually suffer limited quality.
-To remedy this, the last branch of methods (including Mixture-of-Diffusion(Barbero Jiménez 2023), MultiDiffusion (Bar-Tal et al.
-2023), and LLM-grounded Diffusion (Lian et al.
-2023)) first synthesize each object (i.e., bounding box) separately and then compose them to produce the final result.
+To remedy this, the last branch of methods (including Mixture-of-Diffusion[5], MultiDiffusion [4], and LLM-grounded Diffusion [25]) first synthesize each object (i.e., bounding box) separately and then compose them to produce the final result.
 Nevertheless, the computational cost of these methods increases with the number of objects such that the computational overhead is considerable for a text prompt containing plenty of objects.
@@
 Inspired by this, several efforts are made to integrate LLMs to generate layouts from texts, and then utilize layout-conditioned diffusion models for image generation.
-Attention-refocusing (Phung, Ge, and Huang 2024) and LLM-grounded Diffusion (Lian et al.
-2023) explore in-context learning of GPT-4 for layout prediction.
-LayoutGPT (Feng et al.
-2023b) retrieves similar samples from a database as input examples for LLMs.
-LLM-layout-generator (Chen et al.
-2023) leverages Chain-of-Thought (CoT) prompting (Wei et al.
-2022) to improve the layout prediction accuracy.
-However, LLMs are expected to simultaneously perform numerical & spatial reasoning and visual planning of bounding boxes in these methods.
-As a result, inferior accuracy is produced especially for complicated text prompts.
-Contrary to these methods, we propose a divide-and-conquer approach to conduct reasoning and layout prediction separately, which improves the accuracy of layout predictions and substantially advances the prompt fidelity of generated images.
+Attention-refocusing [29] and LLM-grounded Diffusion [25] explore in-context learning of GPT-4 for layout prediction.
+LayoutGPT [14] retrieves similar samples from a database as input examples for LLMs.
+LLM-layout-generator [9] leverages Chain-of-Thought (CoT) prompting [38] to improve the layout prediction accuracy.
+However, these methods all rely on large LLMs such as GPT-4 to directly infer object layouts through end-to-end decoding, which is costly to deploy.
+To the best of our knowledge, using lightweight LLMs for layout prediction in this setting is still under-investigated.
+In this paper, we propose a divide-and-conquer approach that can be applied to lightweight LLMs to enhance the spatial and reasoning accuracy for the layout prediction.
 In addition, during layout-to-image generation, the objects are synthesized progressively according to their difficulties, which achieves not only superior fidelity but also significant computational cost savings.
-Methods Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc. Spatial Layout Image Acc. Acc.
-Stable Diffusion 2.1 - - - - 73.97 53.93 62.38 16.20 - 7.96
-Attend-and-Excite - - - - 76.57 55.94 64.65 18.37 - 9.23
-Divide-and-Bind - - - - 77.95 57.50 66.18 20.15 - 13.15
-Our layout + Layout Guidance - - - - 71.09 45.98 55.84 16.58 - 29.85
-Our layout + GLIGEN - - - - 75.34 62.14 69.01 23.72 - 48.41
-Attention-Refocusing 88.62 91.91 90.44 77.05 77.43 61.78 68.72 25.60 89.57 48.29
-DivCon (Ours) 94.17 89.53 91.80 80.73 78.65 63.41 70.22 29.97 91.87 54.21
-Table 1: Comparison of our method with baselines on HRS benchmark. Text in bold shows the best results of each metric.
-Method
+3 Method
 As illustrated in Fig.
 4, our framework comprises two stages.
-Given a text prompt, LLMs is first adopted to parse it and predict a layout.
+Given a text prompt, a lightweight LLM is first adopted to predict a layout.
 Then, the resultant layout is fed to the diffusion model as an additional condition with the text prompt to achieve text-to-image generation.
-More specifically, each stage is further divided into two steps to handle simple subtasks.
+More specifically, each stage is further divided into two steps to handle subtasks.
 3.1 Text-to-Layout Prediction
-Intuitively, to predict the layout from a text prompt, the number & spatial relations of the objects are first required to be reasoned, and then bounding boxes are generated to describe the coordinates and sizes of the objects.
-However, previous methods commonly conduct these two steps simultaneously without considering their inherent correlation, thereby hindering them from understanding complicated text prompts well.
-To remedy this, we employ the divide-and-conquer strategy to improve the accuracy of predicted layouts.
-Specifically, the text prompt is first fed to an LLM (e.g., GPT-4 (OpenAI 2024)) to perform numerical and spatial reasoning to produce more distinct instructions of objects, as illustrated in Fig.
-4(a).
-Then, these instructions are adopted as additional conditions to predict more accurate bounding boxes.
-(1) Numerical and Spatial Reasoning.
-Given a text prompt $C$ describing an image, an LLM is adopted to predict a set of strings $O=\{s_{j}|j=1,2,...,n\}$ to describe the objects in the prompt.
-If the prompt primarily focuses on counting relationships among objects, each string $s_{j}$ includes both the name and quantity information of each object category $j$.
-If the prompt mainly describes spatial relationships, each string $s_{j}$ contains the object name and its spatial position in the image.
-After numerical and spatial reasoning, the numerical information and spatial relationships can be explicitly parsed from the complicated text prompt as intermedium conditions.
-(2) Bounding Box Prediction.
-After numerical and spatial reasoning, both the meta prompt $C$ and intermedium conditions $O$ are fed to the LLM again to predict a list of name-box pairs $P=\{p_{i}|i=1,2,...,n\}$.
-Each pair $p_{i}=\{o_{i}:(l_{i},t_{i},r_{i},b_{i})|i=1,2,...,n\}$ consists of the object name $o_{i}$ and its bounding box’s coordinates $(l_{i},t_{i},r_{i},b_{i})$.
-Meanwhile, an in-context exemplars including a process of correcting the features of generated bounding boxes is provided to the LLM to improve the quality of the predicted layouts.
+Empirically, we observe that while lightweight LLMs possess basic numerical and spatial reasoning capabilities, these models suffer unstable and inaccurate results for layout generation that requires integrated visual planning.
+To address this, we decouple layout prediction into two explicit stages.
+First, a small-scale and quantized LLM is prompted to extract object-level information from the text prompt, including either numerical counts or spatial positions.
+Then, the same model is prompted to perform visual planning for the layout, guided by the reasoning output through constrained decoding.
+3.1.1 Numerical and Spatial Reasoning.
+Given a text prompt $C$ describing an image, a lightweight LLM is adopted to predict a set of structured strings $O$ containing either numerical or spatial information about objects in the scene.
+If the prompt primarily focuses on counting relationships among objects, the output is parsed into a set of numerical constraints $O_{\text{num}}=\{(c_{j},q_{j})\}_{j=1}^{m}\}$, where $c_{j}$ denotes the object category and $q_{j}$ is its inferred quantity.
+If the prompt mainly describes spatial relationships, the output is parsed into a set of spatial constraints $O_{\text{spa}}=\{(o_{j},p_{j})\}_{j=1}^{n}$, where $o_{j}$ is the object instance and $p_{j}$ is its approximate spatial position (e.g., "top-left").
+These reasoning results are later parsed and converted into token-level constraints that guide layout generation.
+3.1.2 Bounding Box Prediction with Constrained Decoding
+After numerical and spatial reasoning, the meta prompt $C$ are fed to the LLM again with constrains $O$ to predict a list of name-box pairs $P=\{b_{i}|i=1,2,...,n\}$.
+Each pair $b_{i}=\{n_{i}:(x^{(i)}_{1},y^{(i)}_{1},x^{(i)}_{2},y^{(i)}_{2})|i=1,2,...,n\}$ consists of the object name $n_{i}$ and its bounding box’s coordinates.
+To enforce consistency with the predicted constraints, we apply a constrained decoding strategy that inserts structured prefix control tokens during generation.
+For numerical constraints $O_{\text{num}}=\{(c_{j},q_{j})\}_{j=1}^{m}$, we construct a sequence of prefix token sequences:
+$T_{\text{num}}=[\underbrace{\mathtt{c_{1}{:}},\dotsc,\mathtt{c_{1}{:}}}_{q_{1}\ \text{times}},\dotsc,\underbrace{\mathtt{c_{m}{:}},\dotsc,\mathtt{c_{m}{:}}}_{q_{m}\ \text{times}}]$ (1)
+where each prefix token sequence represents a control string (e.g., "apple:") associated with category $c_{j}$, and is inserted before decoding each bounding box.
+This ensures that each category $c_{j}$ appears exactly $q_{j}$ times in the output.
+For spatial constraints $O_{\text{spa}}=\{(o_{j},p_{j})\}_{j=1}^{n}$, we construct a sequence of prefix token sequences:
+$T_{\text{spa}}=[\mathtt{o_{1}(p_{1}):},\mathtt{o_{2}(p_{2}):},\dotsc,\mathtt{o_{n}(p_{n}):}]$ (2)
+where each control prefix string contains both the object name and its position (e.g., "apple(top-left):").
+The token sequence of each prefix is inserted before the decoding of the bounding box for each object $o_{j}$, guiding the predicted location toward the specific region while enabling the model to freely complete the remaining coordinates.
+The detailed prompt templates are provided in the Appendix.
+Method Setting HRS (numerical) HRS (spatial) NSR-1K (numerical) NSR-1K (spatial)
+Prec. Rec. F1. Acc. Acc. Prec. Rec. F1. Acc. Acc.
+TinyLlama 1.1B, FP16 71.80 43.04 53.82 17.87 17.60 90.38 87.46 88.89 70.77 16.25
+TinyLlama + DivCon 1.1B, INT4 81.56 65.14 72.43 31.11 (+13.24) 34.21 (+16.61) 93.45 89.21 91.28 75.99 (+5.22) 34.60 (+18.35)
+Phi-2 2.7B, FP16 76.36 51.43 61.46 23.08 23.56 83.35 95.63 89.07 66.52 39.58
+Phi-2 + DivCon 2.7B, INT4 81.45 53.87 64.85 34.29 (+11.21) 41.51 (+17.95) 87.61 96.05 91.63 80.10 (+13.58) 54.79 (+20.19)
+Mistral-7B 7B, FP16 87.47 67.73 76.35 44.86 28.27 95.39 93.38 94.38 89.06 59.36
+Mistral-7B + DivCon 7B, INT4 93.12 87.45 90.20 69.42 (+24.56) 46.96 (+18.69) 97.36 94.50 95.90 92.03 (+2.97) 69.98 (+10.62)
+Qwen3-4B 4B, FP16 76.56 67.21 71.61 45.15 24.43 89.21 84.07 86.56 76.13 49.66
+Qwen3-4B + DivCon 4B, FP16 88.94 80.95 84.76 65.21 (+20.06) 49.23 (+24.80) 92.21 89.07 91.09 80.21 (+4.08) 59.34 (+9.68)
+Qwen3-8B 8B, INT4 73.27 71.16 72.20 61.16 59.39 91.86 92.04 91.95 86.89 84.45
+Qwen3-8B + DivCon 8B, INT4 93.28 87.76 90.44 74.77 (+13.61) 81.46 (+22.07) 97.45 96.87 97.16 90.07 (+3.18) 94.97 (+10.52)
+GPT-4 (reference) >70B est. 88.13 90.91 89.50 74.05 86.07 97.55 98.41 97.98 93.41 94.68
+Table 1: Quantitative results comparison of our DivCon framework with baselines on HRS and NSR-1K benchmarks in layout prediction.
 3.2 Layout-to-Image Generation
 After layout prediction, the resultant layout is employed as an additional condition to generate the image.
-As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy ones to hard ones in an progressive manner, as shown in Fig.
+As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy ones to hard ones in a progressive manner, as shown in Fig.
 4(b).
@@
 With low-fidelity objects being highlighted, the diffusion model is encouraged to focus on these difficult objects in the second forward pass to achieve higher quality.
-(1) First-Round Generation.
-We employ a layout-to-image diffusion model with sampling optimization (Phung, Ge, and Huang 2024) as the generative model to generate an image conditioned on the text prompt and the predicted layout.
+3.2.1 First-Round Generation.
+We employ a layout-to-image diffusion model with sampling optimization [29] as the generative model to generate an image conditioned on the text prompt and the predicted layout.
 During the denoising process, random noise $z_{T}$ is fed to the diffusion model to gradually denoise it to an image $z_{0}$ after $T$ time steps.
-(2) Consistency Evaluation.
+3.2.2 Consistency Evaluation.
 In the generated image $z_{0}$, each object is cropped using the bounding boxes in the layout and denoted as $x_{i}\in\mathbb{R}^{4\times h\times w}$, where $h\times w$ represents the height and width of the bounding box.
@@
 The higher the similarity is, the higher the fidelity of the corresponding object is.
-Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are ”good enough”.
+Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are "good enough".
 In contrast, other objects with low similarities are considered as hard samples that are more difficult to synthesize (e.g.
-”sheep” and ”horse” in Fig.
+"sheep" and "horse" in Fig.
 3).
 Next, a binary mask $M$ is produced with bounding boxes of easy objects set to 1 and the rest regions set to 0.
-(3) Second-Round Refinement.
-The objective of second-round refinement is to preserve the “good enough” objects while reproducing the remaining ones to ease the difficulty of the diffusion model.
+3.2.3 Second-Round Refinement.
+The objective of second-round refinement is to preserve the “good enough" objects while reproducing the remaining ones to ease the difficulty of the diffusion model.
 To this end, a straightforward approach is to retrieve the latent representation of the objects in the first-round denoising process.
@@
 With the latent representation of high-fidelity objects (masked by $M$) being retrieved, these objects can be well preserved to encourage the model to focus on reconstructing the remaining ones.
-Meanwhile, the bounding boxes of “bad samples” $P^{\prime}$ are passed to the diffusion model as the layout conditions, enabling the model to concentrate more on refining $P^{\prime}$ to further improve the overall quality of the generated image.
-Methods Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc. Spatial Layout Image Acc. Acc.
-Stable Diffusion 2.1 - - - - 78.94 75.42 77.41 33.99 - 22.97
-Attend-and-Excite - - - - 79.21 78.62 78.91 39.51 - 29.61
-Divide-and-Bind - - - - 84.99 79.13 81.96 32.35 - 24.00
-Our layout + Layout Guidance - - - - 81.97 75.41 78.55 43.13 - 52.88
-Our layout + GLIGEN - - - - 85.13 84.97 85.05 52.21 - 59.72
-Attention-Refocusing 98.51 99.11 98.81 94.31 84.61 85.64 85.12 56.10 95.05 69.28
-DivCon (Ours) 98.47 98.35 98.41 94.39 85.41 87.06 86.22 57.35 96.23 73.15
-Table 2: Comparison of our method with baselines on NSR-1K benchmark. Text in bold shows the best results of each metric.
-Experiments
+Meanwhile, the bounding boxes of “bad samples" $P^{\prime}$ are passed to the diffusion model as the layout conditions, enabling the model to concentrate more on refining $P^{\prime}$ to further improve the overall quality of the generated image.
+A more comprehensive analysis of the underlying mechanism is provided in the Appendix.
+Method HRS (Numerical) HRS (Spatial) NSR-1K (Numerical) NSR-1K (Spatial)
+Prec. Rec. F1 Acc. Acc. Prec. Rec. F1 Acc. Acc.
+Stable Diffusion 2.1 73.97 53.93 62.38 16.20 7.96 78.94 75.42 77.41 33.99 22.97
+Attend-and-Excite 76.57 55.94 64.65 18.37 9.23 79.21 78.62 78.91 39.51 29.61
+Divide-and-Bind 77.95 57.50 66.18 20.15 13.15 84.99 79.13 81.96 32.35 24.00
+Our layout + Layout Guidance 71.09 45.98 55.84 16.58 29.85 81.97 75.41 78.55 43.13 52.88
+Our layout + GLIGEN 75.34 62.14 69.01 23.72 42.16 85.13 84.97 85.05 52.21 59.72
+LayoutLLM-T2I 74.38 61.07 67.07 26.56 47.01 83.31 84.01 83.66 55.49 67.99
+Attention-Refocusing 77.43 61.78 68.72 25.60 48.29 84.61 85.64 85.12 56.10 69.28
+DivCon (Ours) 78.65 63.41 70.22 29.97 53.96 85.41 87.06 86.22 57.35 71.65
+Table 2: Quantitative results comparison of our method with baselines on HRS and NSR-1K benchmarks in image generation.
+4 Experiments
 In this section, we provide an evaluation of our method and compare it with state-of-the-art T2I models.
@@
 4.1 Experiment Setup
-(1) Benchmarks and Baselines
-To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS (Bakr et al.
-2023) and NSR-1K (Feng et al.
-2023b).
+4.1.1 Benchmarks and Baselines
+To evaluate the performance of spatial and numerical reasoning in both layouts and images generation tasks, we utilize benchmarks HRS [2] and NSR-1K [14].
 The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships.
@@
 Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations.
-For these benchmarks, Stable Diffusion (Rombach et al.
-2022), Attend-and-Excite (Chefer et al.
-2023), Divide-and-Bind (Li et al.
-2023a), GLIGEN (Li et al.
-2023b), Layout-Guidance (Chen, Laina, and Vedaldi 2023b), and Attention-Refocusing (Phung, Ge, and Huang 2024) are included for comparison.
-In addition, we also include the dataset in (Lian et al.
-2023) for evaluation, which includes four tasks: negation, generative numeracy, attribute binding, and spatial reasoning.
-As our method focuses on numerical and spatial reasoning, we select 100 prompts from these two tasks to construct a benchmark termed LLM-Grounded.
-(2) Evaluation Metrics
-Quality. To assess the quality of generated images, we use FID as the metric (Heusel et al. 2017). Lower FID represents higher similarity between generated images and ground truths.
+To evaluate layout prediction, we employ several open-source LLMs of varying scales and architectures as layout generators to compare their performance, including TinyLlama [45], Phi-2 [19], Mistral-7B [20], Qwen3-4B and Qwen3-8B [37].
+We evaluate both the standalone layout prediction performance of these models and their corresponding quantized versions enhanced by the proposed DivCon framework.
+GPT-4 [28] is included as a reference representing the strongest known performance for LLM’s layout prediction.
+Among them, DivCon with the quantized Qwen3-8B is used to provide layout inputs for the next stage.
+For image generation, Stable Diffusion [34], Attend-and-Excite [7], Divide-and-Bind [23], GLIGEN [24], Layout-Guidance [8], LayoutLLM-T2I [30], and Attention-Refocusing [29] are included for comparison.
+In addition, we also include the dataset in [25] for evaluation, which includes four tasks: negation, generative numeracy, attribute binding, and spatial reasoning.
+As our method focuses on numerical and spatial reasoning, we include all prompts from these two tasks (100 per task) to construct a benchmark termed LLM-Grounded.
+4.1.2 Evaluation Metrics
 Fidelity.
-To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated.
-Particularly, YOLOv8 (Jocher, Chaurasia, and Qiu 2023) is employed to obtain the bounding boxes of the objects in the generated images.
-For counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy (El-Nouby et al.
-2019; Fu et al.
-2020).
-For spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills (Cho, Zala, and Bansal 2022) and LayoutGPT (Feng et al.
-2023b).
+To evaluate the fidelity of both the generated layouts and images, grounding accuracy between layouts/images and the text prompt is calculated.
+YOLOv8 [21] is employed to obtain the bounding boxes of the objects in the generated images.
+For counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy [12, 15].
+For spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills [10] and LayoutGPT [14].
 Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship.
 Otherwise, they are denoted as a top-bottom relationship.
+Quality. To assess the quality of generated images, we use FID as the metric [17]. Lower FID represents higher similarity between generated images and ground truths.
 Efficiency. To validate the efficiency and scalability of our method, inference runtime is evaluated on an AWS g5.xlarge GPU instance.
@@
 4.2 Evaluation Results
-(1) Quantitative Results
-The HRS dataset.
-We present the results on the HRS benchmark in Tab.
+4.2.1 Quantitative Results for Layout Prediction
+We present the results of layout prediction on the HRS and NSR-1K benchmarks in Tab.
 1.
-As we can see, our DivCon outperforms previous T2I models by over 10% in numerical accuracy and about 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance.
-In addition, DivCon surpasses Attention Refocusing (Phung, Ge, and Huang 2024) on most metrics.
-Specifically, our layout prediction improves numerical accuracy by 3% and spatial accuracy by 2%, respectively.
-For image generation, the F1 score and accuracy achieve 2% and 4% gains, while the spatial accuracy is boosted by up to 6%.
-By dividing the T2I task into simple subtasks, our DivCon can better understand the text prompt to achieve higher fidelity.
-The NSR-1K dataset.
-The results on the NSR-1K benchmark are presented in Tab.
+As we can see, all quantized lightweight models, when equipped with our DivCon framework, exhibit substantial improvements across all layout prediction metrics over their original counterparts.
+For instance, Mistral-7B + DivCon achieves a 24.56% gain in numerical accuracy on HRS, while Qwen3-8B + DivCon improves spatial accuracy on HRS by 22.07%.
+This demonstrates the effectiveness of our divide-and-conquer strategy in enhancing reasoning capabilities of lightweight LLMs.
+Notably, our best-performing model, Qwen3-8B + DivCon, surpasses GPT-4 in most HRS’s numerical reasoning metrics, including a 0.94% gain in F1 score and a 0.72% gain in accuracy.
+For HRS’s spatial reasoning, while slightly inferior to GPT-4, it achieves comparable accuracy.
+On NSR-1K, our method matches GPT-4 in both numerical and spatial reasoning.
+These results demonstrate that lightweight open-source models are capable of matching the performance of much larger LLM models.
+4.2.2 Quantitative Results for Image Generation
+We present the results on the HRS and NSR-1K benchmarks in Tab.
 2.
-It can be observed that DivCon outperforms other T2I models by over 20% in numerical accuracy and about 45% in spatial accuracy.
-As compared to Attention Refocusing (Phung, Ge, and Huang 2024), our DivCon produces competitive results in terms of layout accuracy.
-Meanwhile, DivCon achieves approximately 2% gains in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy.
-Quantitative results in terms of image quality are shown in Tab.
+Compared to single-stage text-to-image models, DivCon achieves substantial gains in both numerical and spatial reasoning.
+On HRS, it improves numerical F1 by over 8%, numerical accuracy by 9%, and spatial accuracy by more than 40%.
+On NSR-1K, the corresponding improvements reach 8%, 25%, and 50%, respectively.
+In addition, DivCon also outperforms recent state-of-the-art text-to-layout-to-image models with notable margins.
+On HRS, it yields a 1.5% improvement in numerical F1 and a 3.41% gain in numerical accuracy, while achieving a 5.67% higher spatial accuracy.
+On NSR-1K, DivCon maintains strong performance, surpassing prior methods by 1.1% in numerical F1, 1.25% in accuracy, and 2.37% in spatial accuracy.
+These consistent improvements highlight the effectiveness of our divide-and-conquer strategy in enhancing the fidelity and alignment of generated images under complex reasoning tasks.
+In addition to the HRS and NSR-1K benchmarks, we compare DivCon with LLM-Grounded on numerical and spatial reasoning tasks, as shown in Tab.
 3.
+While both methods achieve comparable performance in numerical reasoning, DivCon demonstrates a notable 5% improvement in spatial accuracy, further highlighting its ability to handle spatial relationships in complex text prompts.
+Methods Numerical Spatial
+LLM-Grounded (GPT4) 84% 82%
+DivCon (Ours) 84% 87%
+Table 3: Comparison with LLM-Grounded
+Image quality results are shown in Tab.
+4.
 As we can see, our DivCon produces comparable FID scores to Stable Diffusion and outperforms attention-refocusing with notable margins.
@@
 Table 4: FID results achieved on NSR-1K.
-LLM-Grounded benchmark.
-Table 4 presents the accuracy achieved by our method and LLM-grounded Diffusion.
-It is clear that our method achieves comparable performance to LLM-Grounded Diffusion on the numerical reasoning while outperforming it by 3% on the spatial reasoning task.
-This also validates the superiority of our method.
-Methods Numerical Spatial
-LLM-Grounded (GPT4) 84% 82%
-DivCon (Ours) 84% 87%
-Table 4: Comparison with LLM-Grounded
-Runtime comparisons are presented in Tab.
+Runtime comparisons of the diffusion process are presented in Tab.
 5.
 Among techniques that require multiple forward passes to synthesize different regions separately, our DivCon produces notable efficiency gains.
-Particularly, LLM-Grounded takes 134.98s to generate an image, while our method achieves a 1.87×1.87$\times$1.87 × speedup.
-Compared with methods with a single forward pass, our method (1 Iter) produces competitive results with comparable efficiency.
+Particularly, LLM-Grounded takes 134.98s to generate an image, while our method achieves a $1.87\times$ speedup.
+Compared with layout-conditioned methods with a single forward pass, our method (1 Iter) produces competitive results with comparable efficiency.
+To conclude, by dividing the T2I task into simpler subtasks and progressively addressing numerical and spatial reasoning, DivCon demonstrates superior fidelity in interpreting text prompts and generating corresponding layouts and images.
+Notably, our framework enables lightweight LLMs to achieve layout prediction performance comparable to large-scale models like GPT-4, significantly reducing computational cost without sacrificing accuracy.
+Moreover, DivCon maintains high visual quality in the generated images while achieving better alignment with the textual descriptions, as reflected in improved spatial and numerical metrics.
+This highlights the robustness and scalability of our approach in tackling challenging T2I scenarios.
 Methods #Forward Pass Runtime
-Stable Diffusion 1 15.81 s
 GLIGEN 1 33.75 s
@@
 Table 5: Comparison of Inference Runtime.
-(2) Qualitative Results
+4.2.3 Qualitative Results
 Figure 5 illustrates the qualitative comparison of DivCon and baselines. In all cases, our DivCon can accurately predict the object quantities and the spatial relations from the text prompts.
@@
 Given a prompt containing more than one object category (Fig.
-5), Stable Diffusion (Rombach et al.
-2022) and Attend-and-Excite (Chefer et al.
-2023) often generate an insufficient quantity of objects, while Attention-Refocusing (Phung, Ge, and Huang 2024) sometimes omits one category of objects or generates an excessive number of objects.
+5), Stable Diffusion [34] and Attend-and-Excite [7] often generate an insufficient quantity of objects, while Attention-Refocusing [29] sometimes omits one category of objects or generates an excessive number of objects.
 For instance, the prompt in Fig.
@@
 In contrast, our DivCon produces accurate object categories and spatial relationships, exhibiting superior performance in these challenging cases.
+Figure 6: User Study: Comparison of image quality.
 User Study.
@@
 6, our method achieves the best performance in terms of average rank, which demonstrates the superior perceptual quality of our results.
-Figure 6: User Study: Comparison of image quality.
 4.3 Ablation Study
@@
 Results are presented in Tab.
-6 and Attention-Refocusing (Phung, Ge, and Huang 2024) is used as the baseline.
+6.
+QWen3-8B [37] and image generator in Attention-Refocusing [29] are used as the baseline for two stages.
 Divide-and-Conquer in Layout Prediction.
 To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step.
-As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 4% in both numerical and spatial accuracy.
+As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 7.59% in numerical F1 score, 8.75% in numerical accuracy and 12.84% in spatial accuracy.
 We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig.
@@
 7 (b)).
-Layout w/ DivCon Image w/ DivCon Numerical Precision Recall F1 Acc. Spatial Acc.
-(1) - - 77.43 61.78 68.72 25.60 48.29
-(2) - 78.26 62.61 69.57 29.46 52.74
-(3) - 78.18 62.36 69.38 27.83 53.17
-(4) 79.13 63.53 70.48 29.62 53.87
-Table 6: Ablation study of our DiVCon on HRS benchmark.
-“Layout w/ DivCon”: Layout prediction with divide-and-conquer strategy.
-“Image w/ DivCon”: Layout-to-Image generation with divide-and-conquer strategy.
-Figure 7: Comparison of the predicted layouts.
-Figure 8: Ablation study of our image generation.
 Divide-and-Conquer in Image Generation.
@@
 As we can see, the divide-and-conquer approach facilitates our method to better reconstruct the objects.
-For example, DivCon accurately reconstructs the small objects (Fig.
-8(a)) and avoids generating the dog outside boxes (Fig.
-8(b)).
-Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy (Tab.
-6).
+For example, as shown in Fig.
+8, the progressive synthesis enabled by the divide-and-conquer strategy allows DivCon to reconstruct objects more accurately, maintaining both numerical consistency and spatial relationships specified in the prompts.
+This demonstrates that the divide-and-conquer framework facilitates better alignment and fidelity in text-to-image tasks.
+Layout w/ DivCon Image w/ DivCon Numerical Precision Recall F1 Acc. Spatial Acc.
+(1) - - 60.15 55.87 57.93 15.96 32.19
+(2) - 76.16 60.31 69.96 23.32 47.06
+(3) - 68.15 59.63 62.27 20.42 41.37
+(4) 78.65 63.41 70.22 29.17 54.21
+Table 6: Ablation study of our DiVCon on HRS benchmark.
+“Layout w/ DivCon": Layout prediction with divide-and-conquer strategy.
+“Image w/ DivCon": Layout-to-Image generation with divide-and-conquer strategy.
+Figure 7: Comparison of the predicted layouts.
+Figure 8: Ablation study of our image generation.
 4.4 Limitations and Discussions
@@
 Figure 9: An illustration of failure case.
-Conclusion
+5 Conclusion
 In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into multiple subtasks to achieve higher image generation quality.
-Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, respectively.
-Then, layout-to-image generation is achieved via an progressive manner to synthesize objects in an easy-to-hard fashion.
+Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, which enables even lightweight LLMs to match the reasoning performance of large-scale models.
+Layout-to-image generation is achieved via a progressive manner to synthesize objects in an easy-to-hard fashion.
 Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.
+Appendix
+In the appendix, we analyze the attention distribution during the second-round forward pass of the diffusion model, and provide example prompts for LLMs.
+Appendix A Attention Distribution in Second-round Generation
+Gated Self-Attention.
+Our model leverages a pretrained layout-to-image generation model (GLIGEN [24]), which utilizes a gated self-attention mechanism to include layout information as new conditions.
+The attention is performed over the concatenation of visual tokens $v$ and grounded layout tokens $h$.
+Since only bad samples are considered during the second-round generation to calculate the attention scores, the model can focus more on these regions.
+To further validate this point, we visualize the attention maps in the last time step.
+As illustrated in Fig.
+10, the attention map obtained in the first round contains high scores outside the groundtruth box.
+In contrast, the attention scores are concentrated in the corresponding box, which ultimately improves the quality of these hard regions.
+Figure 10: The attention map obtained in the first and second round generation.
+Figure 11: Attention maps for word token ‘TV’ in the first and second round generation.
+Number of Numerical Spatial
+Exemplars Pre. Rec. F1. Acc. Acc.
+1 93.25 89.71 91.45 79.14 90.15
+2 94.17 89.63 91.80 80.54 91.87
+3 95.20 87.19 91.02 81.21 91.90
+4 94.56 88.72 91.55 79.21 89.77
+Table 7: Ablation of the number of exemplars on the HRS dataset.
+w/ Negation Numerical Spatial
+Exemplar Pre. Rec. F1. Acc. Acc.
+94.17 89.63 91.80 80.73 91.87
+- 92.15 86.79 89.37 77.24 86.35
+Table 8: Ablation of the negation exemplar on the HRS dataset.
+Cross-Attention Refocusing (CAR) Loss.
+Following [29], our model utilizes a CAR loss for optimization.
+The CAR loss is formulated to enhance the attention scores within the designated box regions while discouraging them outside these areas.
+It is computed individually for each input box and subsequently aggregated across all boxes.
+In our second-round generation, CAR loss is selectively applied to the boxes of hard samples, thereby increasing the relative impact of this loss and improving generation quality in these areas.
+As shown in Fig.
+11, the word token ‘TV’ pays more attention to the area of hard samples instead of being evenly distributed across all samples.
+Appendix B Layout Generation Prompting
+We design prompts for numerical and spatial reasoning respectively, including an in-context exemplar and asks LLM to extract numerical information (e.g.
+5 fire hydrant, 4 cell phone) or location information (e.g.
+dog, left; person, center; cat, center; airplane, right;) of objects in the image (Fig.
+12).
+The bounding box visual planning shares one prompt including 2 in-context exemplars and one negation exemplar which assist the LLM to consider the relative sizes of objects (Fig.
+13).
+Figure 12: The example prompts for numerical and spatial reasoning.
+Figure 13: The example prompts for bounding box visual planning.
+On the HRS benchmark, we conduct ablation experiments of the negation exemplar and the number of in-context exemplars.
+Table 7 compares layout accuracy using varying numbers of in-context exemplars.
+The results show the number of exemplars has minimal impact on the layout accuracy.
+Consequently, we adopt two exemplars in our experiments.
+Table 8 shows that incorporating a negation exemplar yields a 3% improvement in terms of numerical accuracy and a 4% gain on spatial accuracy.
+Consequently, we included the negation exemplar as the default prompt design.